Metrics on State Spaces

This article is dedicated to Richard V. Kadison in anticipation of his completing his seventy-fifth circumnavigation of the sun.

Marc A. Rieffel^{1}

Received: August 17, 1999

Communicated by Joachim Cuntz

Abstract. In contrast to the usual Lipschitz seminorms associated to ordinary metrics on compact spaces, we show by examples that Lipschitz seminorms on possibly non-commutative compact spaces are usually not determined by the restriction of the metric they define on the state space, to the extreme points of the state space. We characterize the Lipschitz norms which are determined by their metric on the whole state space as being those which are lower semicontinuous. We show that their domain of Lipschitz elements can be enlarged so as to form a dual Banach space, which generalizes the situation for ordinary Lipschitz seminorms. We give a characterization of the metrics on state spaces which come from Lipschitz seminorms. The natural (broader) setting for these results is provided by the “function spaces” of Kadison. A variety of methods for constructing Lipschitz seminorms is indicated.

1991 Mathematics Subject Classification: Primary 46L87; Secondary 58B30, 60B10

In non-commutative geometry (based on C^{∗}-algebras), the natural way to specify a metric is by means of a suitable “Lipschitz seminorm”. This idea was first suggested by Connes [C1] and developed further in [C2, C3]. Connes pointed out [C1, C2] that from a Lipschitz seminorm one obtains in a simple way an ordinary metric on the state space of the C^{∗}-algebra. This metric generalizes the Monge–Kantorovich metric on probability measures [KA, Ra, RR]. In this article we make more precise the relationship between metrics on the state space and Lipschitz seminorms. Let ρ be an ordinary metric on a compact space X. The Lipschitz seminorm, L_{ρ}, determined by ρ is defined on functions f on X by

1The research reported here was supported in part by National Science Foundation Grant DMS–96–13833.

(0.1)  L_{ρ}(f) = sup{|f(x) − f(y)|/ρ(x, y) : x ≠ y}.

(This can take value +∞.) It is known that one can recover ρ from L_{ρ} by the relationship

ρ(x, y) = sup{|f(x) − f(y)| : L_{ρ}(f) ≤ 1}.

But a slight extension of this relationship defines a metric, ρ̄, on the space S(X) of probability measures on X, by

(0.2)  ρ̄(μ, ν) = sup{|μ(f) − ν(f)| : L_{ρ}(f) ≤ 1}.

This is the Monge–Kantorovich metric. The topology which it defines on S(X) coincides with the weak-∗ topology on S(X) coming from viewing it as the state space of the C^{∗}-algebra C(X). The extreme points of S(X) are identified with the points of X. On the extreme points, ρ̄ coincides with ρ. Thus the relationship (0.1) can be viewed as saying that L_{ρ} can be recovered just from the restriction of its metric ρ̄ on S(X) to the set of extreme points of S(X). Suppose now that A is a unital C^{∗}-algebra with state space S(A), and let L be a Lipschitz seminorm on A. (Precise definitions are given in Section 2.) Following Connes [C1, C2], we define a metric, ρ, on S(A) by the evident analogue of (0.2). We show by simple finite dimensional examples determined by Dirac operators that L may well not be determined by the restriction of ρ to the extreme points of S(A). It is then natural to ask whether L is determined by ρ on all of S(A), by a formula analogous to (0.1). One of our main theorems (Theorem 4.1) states that the Lipschitz seminorms for which this is true are exactly those which are lower semicontinuous in a suitable sense. For ordinary compact metric spaces (X, ρ) it is known that the space of Lipschitz functions with a norm coming from the Lipschitz seminorm is the dual of a certain other Banach space. Another of our main theorems (Theorem 5.2) states that the same is true in our non-commutative setting, and we give a natural description of this predual. We also characterize the metrics on S(A) which come from Lipschitz seminorms (Theorem 9.11). We should make precise that we ultimately require that our Lipschitz seminorms be such that the metric on S(A) which they determine gives the weak-∗ topology on S(A). An elementary characterization of exactly when this happens was given in [Rf]. (See also [P].) This property obviously holds for finitedimensional C^{∗}-algebras. It is known to hold in many situations for commutative C^{∗}-algebras, as well as for C^{∗}-algebras obtained by combining commutative ones with finite dimensional ones. But this property has not been verified for many examples beyond those. However in [Rf] this property was verified for some interesting infinite-dimensional non-commutative examples, such as the non-commutative tori, and I expect that eventually it will be found to hold in a wide variety of situations. Actually, we will see below that the natural setting for our study is the broader one of order-unit spaces. The theory of these spaces was launched by Kadison in his memoir [K1]. For this reason it is especially appropriate to dedicate this article to him. (In [K2] Kadison uses the terminology “function systems”, but we will follow [Al] in using the terminology “order-unit space” as being a bit more descriptive of these objects.) On the other hand, most of the interesting constructions currently in view of Lipschitz seminorms on non-commutative C^{∗}-algebras, such as those from Dirac operators, or those in [Rf], also provide in a natural way seminorms on all the matrix algebras over the algebras. Thus it is likely that “matrix Lipschitz seminorms” in analogy with the matrix norms of [Ef] will eventually be of importance. But I have not yet seen how to use them in a significant way, and so we do not deal with them here. Let us mention here that a variety of metrics on the state spaces of full matrix algebras have been employed by the practitioners of quantum mechanics. A recent representative paper where many references can be found is [ZS]. We will later make a few comments relating some of these metrics to the considerations of the present paper. The last three sections of this paper will be devoted to a discussion of the great variety of ways in which Lipschitz seminorms can arise, even for commutative algebras. We do discuss here some non-commutative examples, but most of our examples are commutative. I hope in a later paper to discuss and apply some other important classes of non-commutative examples. Some of the applications which I have in mind will require extending the theory developed here to quotients and sub-objects. Finally, we should remark that while we give here considerable attention to how Dirac operators give metrics on state spaces, Connes has shown [C2] that Dirac operators encode far more than just the metric information. In particular they give extensive homological information. But we do not discuss this aspect here. I thank Nik Weaver for suggestions for improvement of the first version of this article, which are acknowledged more specifically below.

1. Recollections on order-unit spaces

We recall [Al] that an order-unit space is a real partially-ordered vector space, A, with a distinguished element e, the order unit, which satisfies:

1) (Order unit property) For each a ∈ A there is an r ∈ ℝ such that a ≤ re.

2) (Archimedean property) If a ∈ A and if a ≤ re for all r ∈ ℝ^{+}, then a ≤ 0.

For any a ∈ A we set

∥a∥ = inf{r ∈ ℝ^{+} : −re ≤ a ≤ re}.

We obtain in this way a norm on A. In turn, the order can be recovered from the norm, because 0 ≤ a ≤ e iff ∥a∥ ≤ 1 and ∥e − a∥ ≤ 1. The primary source of examples consists of the linear spaces of all self-adjoint elements in unital C^{∗}-algebras, with the identity element serving as order unit. But any linear space of bounded self-adjoint operators on a Hilbert space will be an order-unit space if it contains the identity operator. We expect that this broader class of examples will be important for the applications of metrics on state spaces. We will not assume that A is complete for its norm. This is important for us because the domains of Lipschitz norms will be dense, but usually not closed, in the completion. (The completion is always again an order-unit space.) This also accords with the definition in [Al]. By a state of an order-unit space (A, e) we mean a continuous linear functional, μ, on A such that μ(e) = 1 = ∥μ∥. States are automatically positive. We denote the collection of all the states of A, i.e. the state space of A, by S(A). It is a w^{∗}-compact convex subset of the Banach space dual, A′, of A. To each a ∈ A we get a function, â, on S(A) defined by â(μ) = μ(a). Then â is an affine function on S(A) which is continuous by the definition of the w^{∗}-topology. The basic representation theorem of Kadison [K1, K2, K3] (see Theorem II.1.8 of [Al]) says that for any order-unit space the representation a → â is an isometric order isomorphism of A onto a dense subspace of the space Af(S(A)) of all continuous affine functions on S(A), equipped with the supremem norm and the usual order on functions (and with e clearly carried to the constant function 1). In particular, if A is complete, then it is isomorphic to all of Af(S(A)). Thus we can view the order-unit spaces as exactly the dense subspaces containing 1 inside Af(K), where K is any compact convex subset of a topological vector space. This provides an effective view from which to see many of the properties of order-unit spaces. Most of our theoretical discussion will be carried out in the setting of order-unit spaces and Af(K), though our examples will usually involve specific C^{∗}-algebras. We let C(K) denote the real C^{∗}-algebra of all continuous functions on K, in which Af(K) sits as a closed subspace. It will be important for us to work on the quotient vector space Ã = A/(ℝe). We let ∥ ∥^{∼} denote the quotient norm on Ã from ∥ ∥. This quotient norm is easily described. For a ∈ A set

max(a) = inf{r : a ≤ re}
min(a) = sup{r : re ≤ a},

so that ∥a∥ = (max(a)) ∨ (− min(a)). Then it is easily seen that

∥ã∥^{∼} = (max(a) − min(a))/2.

2. The radius of the state space

Let A be an order-unit space. Since the term “Lipschitz seminorm” has somewhat wide but imprecise usage, we will not use this term for our main objects of precise study (which we will define in Section 5). Almost the minimal requirement for a Lipschitz seminorm is that its null-space be exactly the scalar multiples of the order unit. We will use the term “Lipschitz seminorm” in this general sense. We emphasize that a Lipschitz seminorm will usually not be continuous for ∥ ∥. Let L be a Lipschitz seminorm on A. For μ, ν ∈ S(A) we can define a metric, ρ_{L}, on S(A) by

ρ_{L}(μ, ν) = sup{|μ(a) − ν(a)| : L(a) ≤ 1}

(which may be +∞). Then ρ_{L} determines a topology on S(A). Eventually we want to require that this topology agrees with the weak-∗ topology. Since S(A) is weak-∗ compact, ρ_{L} must then give S(A) finite diameter. We examine this latter aspect here, in part to establish further notation. It is actually more convenient for us to work with “radius” (half the diameter), since this will avoid factors of 2 in various places. We would like to use the properties of order-unit spaces to express the radius in terms of L in a somewhat more precise way than was implicit in [Rf] in its more general context. The following considerations [Al] will also be used extensively later. As in [Rf] and in the previous section, we denote the quotient vector space A/(ℝe) by Ã, with its quotient norm ∥ ∥^{∼}. But in addition to this norm, the quotient seminorm L̃ from L is also a norm on Ã, since L takes value 0 only on ℝe. The dual Banach space to Ã for ∥ ∥^{∼} is just A′^{0}, the subspace of A′ consisting of those λ ∈ A′ such that λ(e) = 0. We denote the norm on A′ dual to ∥ ∥ still by ∥ ∥. The dual norm on A′^{0} is just the restriction of ∥ ∥ to A′^{0}. If we view A as a dense subspace of Af(K) ⊆ C(K), then by the Hahn–Banach theorem λ extends (not uniquely) to C(K) with same norm. There we can take the Jordan decomposition into disjoint non-negative measures. Note that for positive measures their norm on C(K) equals their norm on A, since e ∈ A. Thus we find μ, ν ≥ 0 such that λ = μ − ν and ∥λ∥ = ∥μ∥ + ∥ν∥. But 0 = λ(e) = μ(e) − ν(e) = ∥μ∥ − ∥ν∥. Consequently ∥μ∥ = ∥ν∥ = ∥λ∥/2. Thus if ∥λ∥ ≤ 2 we have ∥μ∥ = ∥ν∥ ≤ 1. If ∥λ∥ < 2 set t = ∥μ∥ < 1, and rescale μ and ν so that they are in S(A). Then

λ = tμ − tν = μ − (tν + (1 − t)μ).

Now (tν + (1 − t)μ) is no longer disjoint from μ, but we have obtained the following lemma, which will be used in a number of places.

2.1 Lemma. The ball D_{2} of radius 2 about 0 in A′^{0} coincides with {μ − ν : μ, ν ∈ S(A)}.

Notice that if there is an a ∈ A such that L(a) = 0 but a ∉ ℝe, then from this lemma we can find μ, ν ∈ S(A) such that (μ−ν)(a) ≠ 0, so that ρ_{L}(μ, ν) = +∞.

Thus our standing assumption that there is no such a serves to reduce the possibility of having infinite distances. But it does not eliminate this possibility, as seen by the example of the algebra of smooth (or Lipschitz) functions of compact support on the real line, with constant functions adjoined, and with the usual Lipschitz seminorm.

2.2 Proposition. With notation as earlier, the following conditions are equivalent for an r ∈ ℝ^{+}:

1) For all μ, ν ∈ S(A) we have ρ_{L}(μ, ν) ≤ 2r.

2) For all a ∈ A we have ∥ã∥^{∼} ≤ rL^{∼}(ã).

Proof. Suppose that condition 1 holds. Let a ∈ A and λ ∈ D_{2}. Then by the lemma λ = μ − ν for some μ, ν ∈ S(A). Thus

|λ(a)| = |(μ − ν)(a)| ≤ L(a)ρ_{L}(μ, ν) ≤ L(a)2r.

Since λ(e) = 0, thus inequality holds whenever a is replaced by a+se for s ∈ ℝ. Thus condition 2 holds. Conversely, suppose that condition 2 holds. Then for any μ, ν ∈ S(A) and a ∈ A with L(a) ≤ 1 we have

|μ(a) − ν(a)| = |(μ − ν)(a)| ≤ 2∥ã∥^{∼} ≤ 2r.

Thus ρ_{L}(μ, ν) ≤ 2r as desired. □

Of course, we call the smallest r for which the conditions of this proposition hold the radius of S(A). We caution that just because a metric space has radius r, it does not follow that there is a ball of radius r which contains it, as can be seen by considering equilateral triangles in the plane. We remark that just because ρ_{L} gives S(A) finite radius, it does not follow that ρ_{L} gives the weak-∗ topology. Perhaps the simplest example arises when A is infinite dimensional and L(a) = ∥ã∥^{∼}.

3. Lower semicontinuity for Lipschitz seminorms

Let L be any Lipschitz seminorm on an order-unit space A. (We will not at first require that it give S(A) finite diameter.) We would like to show that L and ρ_{L} contain the same information, and more specifically that we can recover L from ρ_{L} as being the usual Lipschitz seminorm for ρ_{L}. By this we mean the following. Let ρ be any metric on S(A), possibly taking value +∞. Define L_{ρ} on C(S(A)) by

(3.1)  L_{ρ}(f) = sup{|f(μ) − f(ν)|/ρ(μ, ν) : μ ≠ ν},

where this may take value +∞. Let Lip_{ρ} = {f : L_{ρ}(f) < ∞}. We can restrict L_{ρ} to Af(S(A)). In general, few elements of Af(S(A)) will be in Lip_{ρ}. However, on viewing the elements of A as elements of Af(S(A)), we have:

3.2 Lemma. Let L be a Lipschitz seminorm on A with corresponding metric ρ_{L} on S(A). Then A ⊆ Lip_{ρ_{L}}, and on A we have L_{ρ_{L}} ≤ L , in the sense that L_{ρ_{L}}(a) ≤ L(a) for all a ∈ A.

Proof. For μ, ν ∈ S(A) and a ∈ A we have

|â(μ) − â(ν)| = |μ(a) − ν(a)| ≤ L(a)ρ_{L}(μ, ν).

□

For later use we remark that if L and M are Lipschitz seminorms on A and if M ≤ L, then ρ_{M} ≥ ρ_{L} in the evident sense. We would like to recover L on A from ρ_{L} by means of formula (3.1). However, the seminorms defined by (3.1) have an important continuity property:

3.3 Definition. Let A be a normed vector space, and let L be a seminorm on A, except that we permit it to take value +∞. Then L is lower semicontinuous if for any sequence {a_{n}} in A which converges in norm to a ∈ A we have L(a) ≤ lim inf{L(a_{n})}. Equivalently, for one, hence every, t ∈ ℝ with t > 0, the set

L_{t} = {a ∈ A : L(a) ≤ t}

is norm-closed in A.

3.4 Proposition. Let A be an order-unit space, and let ρ be any metric on S(A), possibly taking value +∞. Define L_{ρ} on C(S(A)) by formula (3.1). Then L_{ρ} is lower semicontinuous. Consequently, the restriction of L_{ρ} to any subspace of C(S(A)), such as A or Af(S(A)), will be lower semicontinuous.

Proof. When we view L_{ρ} as a function of f, the formula (3.1) says that L_{ρ} is the pointwise supremum of a collection of functions (labeled by pairs μ, ν with μ ≠ ν) which are clearly continuous on C(S(A)) for the supremum norm. But the pointwise supremum of continuous functions is lower semicontinuous. □

3.5 Example. Here is an example of a Lipschitz seminorm L whose metric can be seen to give S(A) the weak-∗ topology, but which is not lower semicontinuous. Let I = [−1, 1], and let A = C^{1}(I), the algebra of functions which have continuous first derivatives on I. Define L on A by

L(f) = ∥f′∥_{∞} + |f′(0)|.

For each n let g_{n} be the function defined by g_{n}(t) = n|t| for |t| ≤ 1/n, and g_{n}(t) = 1 elsewhere. Let f_{n}(t) = ∫_{−1}^{t} g_{n}(s)ds. Then the sequence {f_{n}} converges uniformly to the function f given by f(t) = t + 1. But L(f_{n}) = 1 for each n, whereas L(f) = 2.

A substantial supply of examples of lower semicontinuous seminorms can be obtained from the W^{∗}-derivations of Weaver [W2, W3]. These derivations will in general have large null spaces, and the seminorms from them need not give the weak-∗ topology on the state space. But many of the specific examples of W^{∗}-derivations which Weaver considers do in fact give the weak-∗ topology. In terms of Weaver’s terminology, which we do not review here, we have:

3.6 Proposition. Let M be a von Neumann algebra and let E be a normal dual operator M-bimodule. Let δ : M → E be a W^{∗}-derivation, and denote the domain of δ by L, so that L is an ultra-weakly dense unital ∗-subalgebra of M. Define a seminorm, L, on L by L(a) = ∥δ(a)∥_{E}. Then L is lower semicontinuous, and L_{1} = {a ∈ L : L(a) ≤ 1} is norm-closed in M itself.

Proof. Let {a_{n}} be a sequence in L which converges in norm to b ∈ M. To show that L is lower semicontinuous, it suffices to consider the case in which {a_{n}} is contained in L_{1}. Then the set {(a_{n}, δ(a_{n}))} is a bounded subset of the graph of δ for the norm max{∥ ∥_{M}, ∥ ∥_{E}}. Since the graph of a W^{∗}-derivation is required to be ultra-weakly closed, and since bounded ultraweakly closed subsets are compact for the ultra-weak topology, there is a subnet which converges ultra-weakly to an element (c, δ(c)) of the graph of δ. Then necessarily c = b, so that b ∈ L, and δ(b) is in the ultra-weak closure of {δ(a_{n})}. Consequently L(b) = ∥δ(b)∥ ≤ 1. □

Because of the importance of Dirac operators, it is appropriate to verify lower semicontinuity for the Lipschitz seminorms which they determine. This is close to a special case of Proposition 3.6, but does not require any kind of completeness, nor an algebra structure on A.

3.7 Proposition. Let A be a linear subspace of bounded self-adjoint operators on a Hilbert space H, containing the identity operator. Let D be an essentially self-adjoint operator on H whose domain, D(D), is carried into itself by each element of A. Assume that [D, a] is a bounded operator on D(D) for each a ∈ A (so that [D, a] extends uniquely to a bounded operator on H). Define L on A by L(a) = ∥[D, a]∥. Then L is lower semicontinuous.

Proof. Let {a_{n}} be a sequence in A which converges in norm to a ∈ A. Suppose that there is a constant, k, such that L(a_{n}) ≤ k for all n. For any ξ, η ∈ D(D) with ∥ξ∥ = 1 = ∥η∥ we have

⟨[D, a]ξ, η⟩ = ⟨aξ, Dη⟩ − ⟨Dξ, aη⟩ = lim⟨[D, a_{n}]ξ, η⟩.

But |⟨[D, a_{n}]ξ, η⟩| ≤ k for each n, and so ∥[D, a]∥ ≤ k. □

We remark that the Lipschitz seminorms constructed in [Rf] by means of actions of compact groups are easily seen to be lower semicontinuous.

4. Recovering L from ρ_{L}

In this section we show that a lower semicontinuous Lipschitz seminorm L can be recovered from its metric ρ_{L}. But before showing this we would like to emphasize the following point. Let (X, ρ) be an ordinary compact metric space, with A the algebra of its Lipschitz functions, with Lipschitz seminorm L. Then S(A) consists of the probability measures on X, and the points of X correspond exactly to the extreme points of S(A). The restriction of ρ_{L} to the extreme points is exactly ρ. Thus when one says that one can recover L from the metric ρ, one is saying that one can recover L from the restriction of ρ_{L} on S(A) to the extreme points of S(A). However, for the more general situation which we are considering, it will be false in general that we can recover L from the restriction of ρ_{L} to the extreme points of S(A). Simple explicit examples will be given in Section 7. One of the main theorems of this paper is:

4.1 Theorem. Let L be a lower semicontinuous Lipschitz seminorm on an order-unit space A, and let ρ_{L} denote the corresponding metric on S(A), possibly taking value +∞. Let L_{ρ_{L}} be defined by formula (3.1), but restricted to A ⊆ Af(S(A)). Then

L_{ρ_{L}} = L.

Theorem 4.1 is an immediate consequence of the following theorem, since we saw that lower semicontinuity coincides with L_{1} being norm closed.

4.2 Theorem. Let L be any Lipschitz seminorm on an order-unit space A, and let ρ_{L} denote the corresponding metric on S(A). Let L_{ρ_{L}} be defined by formula (3.1), but restricted to A ⊆ Af(S(A)). Then {a ∈ A : L_{ρ_{L}}(a) ≤ 1} coincides with the norm closure, L̄_{1}, of L_{1} in A. In particular, L_{ρ_{L}} is the largest lower semicontinuous seminorm smaller than L, and ρ_{L_{ρ_{L}}} = ρ_{L}.

Proof. (An idea leading to this proof, which is simpler than my original proof, was suggested to me by Nik Weaver.) On A′ we define the seminorm, L′ , dual to L, by

L′(λ) = sup{|λ(a)| : L(a) ≤ 1}.

Note that L′ takes value +∞ on any λ for which λ(e) ≠ 0, and very possibly on some elements of A′^{0} as well. But at any rate we have the following key relationship:

4.3 Lemma. For μ, ν ∈ S(A) we have ρ_{L}(μ, ν) = L′(μ − ν).

Proof.

L′(μ − ν) = sup{|(μ − ν)(a)| : L(a) ≤ 1}
= sup{|μ(a) − ν(a)| : L(a) ≤ 1} = ρ_{L}(μ, ν).

□

Because L_{1} is already convex and balanced, the bipolar theorem [Cw] says that L̄_{1} is exactly the bipolar of L_{1}. Thus we just need to show that {a ∈ A : L_{ρ_{L}}(a) ≤ 1} is the bipolar of L_{1}. Now it is clear that the unit L′-ball in A′ is exactly the polar [Cw] of L_{1}. This provides the last of the following equivalences. Let a ∈ A. Then: L_{ρ_{L}}(a) ≤ 1 exactly if |μ(a) − ν(a)| ≤ ρ_{L}(μ, ν) for all μ, ν ∈ S(A) , exactly if |λ(a)| ≤ L′(λ) for all λ ∈ D_{2} (by Lemma 4.3 and Lemma 2.1), exactly if |λ(a)| ≤ 1 for all λ ∈ A′ with L′(λ) ≤ 1, exactly if a is in the prepolar of {λ : L′(λ) ≤ 1} (by definition [Cw]), exactly if a is in the bipolar of L_{1}. It is clear that L_{ρ_{L}} is lower semicontinuous, that it is the largest such seminorm smaller than L, and that it gives the same metric. □

Note in particular that if L gives S(A) finite diameter, or the weak-∗ topology, then so does L_{ρ_{L}}. We remark that a sort of dual version of Theorem 4.1 can be found later in Theorem 9.7. We have the following related considerations. Suppose again that L is a Lipschitz seminorm on an order-unit space A. Let Ā denote the completion of A for ∥ ∥, and let L̄_{1} denote now the closure of L_{1} in Ā rather than just in A. Let L̄ denote the corresponding “Minkowski functional” on Ā obtained by setting, for b ∈ Ā,

L̄(b) = inf{r ∈ ℝ^{+} : b ∈ rL̄_{1}}.

Since there may be no such r, we must allow the value +∞. With this understanding, L̄ will be a seminorm on Ā. It is easily seen that L̄(b) ≤ 1 exactly if b ∈ L̄_{1}, and that L̄ is lower semicontinuous because L̄_{1} is closed. Up to this point we did not require lower semicontinuity of L. It’s import is given by:

4.4 Proposition. Let L be a lower semicontinuous Lipschitz seminorm on an order-unit space A. Let L̄ on Ā be defined as above. Then L̄ is an extension of L, that is, for a ∈ A we have L̄(a) = L(a). Furthermore, ρL̄ = ρ_{L}.

Proof. Suppose that a ∈ A and L(a) = 1. Then a ∈ L_{1} ⊆ L̄_{1} and so clearly L̄(a) ≤ 1. Conversely, if L̄(a) ≤ 1, then a ∈ L̄_{1}. Thus there is a sequence {a_{n}} in L_{1} which converges to a, with L(a_{n}) ≤ 1 for every n. From the lower semicontinuity of L it follows that L(a) ≤ 1. Finally, for μ, ν ∈ S(A) we have

ρL̄(μ, ν) = sup{|μ(a)−ν(a)| : a ∈ L̄_{1}} = sup{|μ(a)−ν(a)| : a ∈ L_{1}} = ρ_{L}(μ, ν).

□

Note in particular that if L gives S(A) finite diameter, or the weak-∗ topology, then so does L̄. However, in general L̄ need not be a Lipschitz seminorm. For example, let A be the algebra of real polynomials viewed as functions on the interval [0, 2], and let L be the usual Lipschitz seminorm but defined using only points in [0, 1].

4.5 Definition. We will call L̄ the closure of L. We will say that a Lipschitz seminorm is closed if L = L̄ (on the subspace where L̄ is finite), or equivalently, if L_{1} is complete for the metric from ∥ ∥.

Then Proposition 4.4 says that for most purposes we can assume that L is closed if convenient.

Suppose now that L is a Lipschitz seminorm on A which is closed. On A we can define a new norm, |∥ ∥|, by

|∥a∥| = ∥a∥ + L(a).

It is easily verified that A is complete for this new norm. Suppose that A is a ∗-algebra and ∥ ∥ is a C^{∗}-norm (this can be weakened). Suppose further that L is a closed Lipschitz seminorm on A which satisfies the Leibniz inequality. Then the new norm is a normed-algebra norm, and so A becomes a Banach algebra for the new norm. In Sections 10 and 11 we will indicate many examples of Lipschitz seminorms satisfying the Leibniz inequality. This provides a rich class of examples of Banach algebras which merit study (even in the cases when they are commutative) along the lines considered in [BCD, J, W1].

5. The pre-dual of (Ã, L̃)

It has been shown in an increasing variety of situations that the space of Lipschitz functions with a suitable Lipschitz norm is isometrically isomorphic to the dual of some Banach space. Some of the history of this phenomenon is sketched in the notes at the end of chapter 2 of [W1], or more briefly in [W2]. Within the non-commutative setting, Weaver shows in Proposition 2 of [W2] that the domains of W^{∗}-derivations (as defined there) are dual spaces. However, his W^{∗}-derivations can have large null spaces, and they need not give the weak-∗ topology on S(A). Nevertheless, Weaver’s approach applies to the non-commutative tori, and gives them the same space of Lipschitz elements as the approach of the present paper (when combined with [Rf]). In fact, Weaver shows in [W3] that for the non-commutative tori one can also define Lip^{α}, and that Lip^{α} is actually the second dual of lip^{α} when α < 1. To show within our setting that the space of Lipschitz elements is the dual of a Banach space, we need to assume that ρ_{L} gives the weak-∗ topology on S(A). As before, let L_{1} = {a : L(a) ≤ 1}. From theorem 1.8 of [Rf] we know that ρ_{L} will give the weak-∗ topology on S(A) exactly if the image of L_{1} in Ã is totally bounded for ∥ ∥^{∼}. Equivalently, by theorem 1.9 of [Rf], L must give S(A) finite radius, and for one, hence all, t ∈ ℝ with t > 0, the set

B_{t} = {a : L(a) ≤ 1 and ∥a∥ ≤ t}

must be totally bounded in A for ∥ ∥. We remark that this implies that if {a_{n}} is a sequence (or net) in A converging pointwise on S(A) to a ∈ A, and if there is a constant k such that ∥a_{n}∥ ≤ k and L(a_{n}) ≤ k for all n, then a_{n} converges to a in norm. This is because {a_{n}} is contained in kB_{1} whose closure in the completion Ā of A is compact. Let b be any norm limit point of {a_{n}} in Ā. Then a subsequence of {a_{n}} converges in norm to b. But it still converges pointwise on S(A) to a. Consequently b = a, and a is the only norm limit point of {a_{n}}. We now have in view all the requirements on Lipschitz seminorms which we need for our present purposes. So we now define what we expect is the correct way to specify metrics on compact non-commutative spaces:

5.1 Definition. Let A be an order-unit space. By a Lip-norm on A we mean a seminorm, L, on A (taking finite values) with the following properties:

1) For a ∈ A we have L(a) = 0 if and only if a ∈ ℝe.

2) L is lower semicontinuous.

3) {a ∈ A : L(a) ≤ 1} has image in Ã which is totally bounded for ∥ ∥^{∼}.

We remark that it is easily checked that the closure (Definition 4.5) of a Lipnorm is again a Lip-norm. Within the present setting the fact that the space of Lipschitz elements is a dual Banach space takes the following form (which requires the Lip-norm to be closed).

5.2 Theorem. Let A be an order-unit space, and let L be a Lip-norm on A which is closed. Let K = {ã ∈ Ã : L̃(ã) ≤ 1}, so that K is a compact (convex) set for ∥ ∥^{∼}. Then (Ã, L̃) is naturally isometrically isomorphic to the dual Banach space of Af_{0}(K), the Banach space of continuous affine functions on K which take value 0 at 0 ∈ Ã, with the supremum norm.

Proof. Let L_{1} and B_{t} be as defined as above. Because L is closed, the totally bounded sets B_{t} are complete for ∥ ∥, and so are compact. From the finite radius considerations of Section 2 the image of L_{1} in Ã will coincide with the image of B_{t} for sufficiently large t. Hence the image of L_{1} in Ã is compact for ∥ ∥^{∼}, not just totally bounded. But the image of L_{1} is exactly K as defined in the statement of the theorem. We can now argue as in the proof of proposition 1 of [W4]. We include the argument here in a form specific to our particular situation. Let V = Af_{0}(K), as defined in the statement of the theorem. Then from lemma 4.1 of [K3] each element of V extends to a linear functional (not necessarily continuous for ∥ ∥^{∼}) on Ã. But we still view V as equipped with the uniform norm ∥ ∥_{∞} from C(K), for which V is complete. Then for any f ∈ V we have

∥f∥_{∞} = sup{f(ã) : ã ∈ K} = sup{f(ã) : L̃(ã) ≤ 1}.

Consequently ∥ ∥_{∞} is just the dual norm to the norm L̃ on Ã. But V will usually be much smaller than the entire dual Banach space of (Ã, L̃) because of the requirement that if f ∈ V then f is continuous on K. We let V′ denote the dual Banach space to V. We have the evident mapping σ from Ã to V′ defined by σ(ã)(f) = f(ã). Use of the Hahn–Banach theorem shows that Af_{0}(K) separates the points of K, and from this we see that σ is injective. Furthermore |σ(ã)(f)| = |f(ã)| ≤ ∥f∥_{∞}L̃(ã), and so ∥σ∥ ≤ 1 for the norm L̃ on Ã. In particular, σ(K) ⊆ (V′)_{1}, the unit ball of V′. From the definitions of σ and V we see immediately that σ is continuous from K to (V′)_{1} with its weak-∗ topology from V. Since K is compact, σ(K) must be compact for the weak-∗ topology. If σ(K) were not all of (V′)_{1}, there would be a φ_{0} ∈ (V′)_{1} and a weak-∗ continuous linear functional separating φ_{0} from σ(K). But every weak-∗ linear functional comes from V. Thus there would be an f ∈ V such that

f(ã) ≤ 1 < φ_{0}(f)

for every ã ∈ K. But the first inequality means that ∥f∥_{∞} ≤ 1, and so the second inequality means that ∥φ_{0}∥ > 1, contradicting the assumption that φ_{0} ∈ (V′)_{1}. Thus σ(K) = (V′)_{1}. Consequently σ is an isometric isomorphism of (Ã, L̃) with V′. □

We remark that, if desired, we can make A itself into the dual of a Banach space, in a non-canonical way, as follows. Let r be the radius of (A, L), and let μ be any fixed state of A. Define an actual norm, L_{μ}, on A by

L_{μ}(a) = max{|μ(a)|/r, L(a)}.

Let L̃_{μ} be the quotient of L_{μ} on Ã. It is clear that L̃_{μ} ≥ L̃. But for any given a ∈ A we can find α ∈ ℝ such that ∥a−α∥ ≤ rL̃(ã), by the definition of radius. Then

|μ(a − α)| ≤ ∥a − α∥ ≤ rL̃(ã),

while L(a − α) = L̃(ã). Consequently L̃_{μ}(ã) ≤ L̃(ã), so that, in fact, L̃_{μ} = L̃. Thus (A, L_{μ}) has (Ã, L̃) as quotient space. The quotient map splits by the isometric map ã ↦ a − μ(a). Since (Ã, L̃) is isometrically isomorphic to a dual Banach space, it follows easily that (A, L_{μ}) is also. See also section 2 of [H], which gives a slightly different approach because the norm on Lip_{ρ} is slightly different from that implicit here. Let K and V = Af_{0}(K) be as in the statement of Theorem 5.2. As in Section 2, the dual of (Ã, ∥ ∥^{∼}) is A′^{0}. By the finite diameter condition and Proposition 2.2 each λ ∈ A′^{0} defines a continuous linear functional on (Ã, L̃). Each such functional is clearly continuous on K for its topology from ∥ ∥^{∼}. Thus each λ ∈ A′^{0} defines an element of V, and so we obtain a linear map from A′^{0} into V. From Theorem 5.2 the norm ∥ ∥_{∞} on V from C(K) coincides with the dual norm L′ from (Ã, L̃). We have the following addition to Theorem 5.2.

5.3 Proposition. The image of A′^{0} in Af_{0}(K) is dense in Af_{0}(K) for its norm ∥ ∥_{∞} = L′.

Proof. Let φ be any continuous linear functional on V which is 0 on the image of A′^{0}. From Theorem 5.2 every continuous linear functional on V comes from an element of Ã. If ã is the element of Ã corresponding to φ, we then have λ(ã) = 0 for all λ ∈ A′^{0}, which implies that ã = 0 so that φ ≡ 0. It follows from the Hahn–Banach theorem that the image of A′^{0} is norm dense in V. □

6. Extreme points

Let L be a Lipschitz seminorm on an order-unit space A, and let ρ_{L} be the corresponding metric on S(A). Let E denote the set of extreme points of S(A).

Then E need not be a closed subset of S(A), but S(A) is the closed convex hull of E by the Krein–Milman theorem. Of course ρ_{L} restricts to a metric on E. We will give explicit examples in the next section to show that even when L is a Lip-norm the restriction of ρ_{L} to E does not determine ρ_{L} or L. Nevertheless, we can try to use the restriction of ρ_{L} to define a new Lipschitz seminorm, L^{e}, on A, by

L^{e}(a) = sup{|ε(a) − η(a)|/ρ_{L}(ε, η) : ε, η ∈ E, ε ≠ η}.

6.1 Proposition. With the above definition, L^{e} is a lower semicontinuous Lipschitz seminorm on A, and it is the smallest such on A whose metric on S(A) agrees on E with that of L. If L is a Lip-norm then so is L^{e}.

Proof. From Theorem 4.2 it is clear that we can assume that L is lower semicontinuous. From Theorem 4.1 we know that any lower semicontinuous Lipschitz seminorm, say L_{1}, is recovered from its metric by a supremum as above, but ranging over all of S(A) rather than just over E. Thus if the metric for L_{1} agrees with ρ_{L} on E, we must have L^{e} ≤ L_{1}. By using the argument in the proof of Proposition 3.4 it is easily seen that L^{e} is lower semicontinuous. Suppose that L^{e}(a) = 0 for some a ∈ A. Recall that D_{2} = {λ ∈ A′^{0} : ∥λ∥ ≤ 2}.

6.2 Lemma. The convex hull of {ε − η : ε, η ∈ E, ε ≠ η} is dense in D_{2} for the weak-∗ topology.

Proof. From Lemma 2.1 we know that any element of D_{2} can be expressed as μ − ν for μ, ν ∈ S(A). By the Krein–Milman theorem each of μ, ν can be approximated arbitrary closely in the weak-∗ topology by convex combinations from E, say ∑ α_{j}ε_{j} and ∑ β_{k}η_{k}. But the difference of such combinations can be expressed as

∑(α_{j}β_{k})(ε_{j} − η_{k}).

□

From this lemma it is clear that if L^{e}(a) = 0 then L(a) = 0, and thus a ∈ ℝe. Also, it is easy to see that ρ_{L^{e}} agrees with ρ_{L} on E. Finally, we must show that if L is a Lip-norm then the image of K_{0} = {a : L^{e}(a) ≤ 1} in Ã is totally bounded for ∥ ∥^{∼}. Notice that this image is larger than that for L, so we can not immediately apply the corresponding fact for L. Let Ē denote the closure of E in S(A). It is clear that the supremum defining L^{e} could just as well be taken over Ē, and so L^{e} on A is just the Lipschitz norm for the metric ρ_{L} restricted to Ē. Thus K_{0} can be viewed as contained in {f ∈ C(Ē) : L^{e}(f) ≤ 1}, and the latter has totally bounded image in C(Ē)/ℝe since it consists of Lipschitz functions for a metric and Ē is compact. Thus K_{0} has totally bounded image in C(Ē)/ℝe. But the restriction map from Af(S(A)) to C(Ē) is isometric for ∥ ∥_{∞} since Ē contains the extreme points. (See Theorem II.1.8 of [Al]. We are dealing here with Kadison’s smallest separating representation.) It follows easily that K_{0} has totally bounded image in Ã as needed. □

We remark that if F is any subset of S(A) which contains E, then we can use F instead of E to define a Lip-norm L^{F} just as we defined L^{e} above. Then we will have

L^{e} ≤ L^{F} ≤ L

in the evident sense, with reverse inequalities for the corresponding metrics. Suppose that A is a dense ∗-subalgebra of a C^{∗}-algebra, Ā, and that L is a Lipnorm on A, with corresponding metric ρ_{L} on S(A). As above let E denote the set of extreme points of S(A). Assume first that A is commutative. Then E is compact and Ā ≅ C(E). Assume that L = L^{e}. Then L is the usual Lipschitz norm coming from the metric on the compact set E obtained by restricting ρ_{L} to E. But in this case we know that L must then satisfy the Leibniz rule

L(ab) ≤ L(a)∥b∥ + ∥a∥L(b).

It is thus natural to ask the general question:

6.3 Question. What conditions on a Lip-norm L on a general unital C^{∗}- algebra imply that L satisfies the Leibniz rule?

In the next section we will see examples of Lip-norms which do not satisfy L = L^{e} and yet satisfy the Leibniz rule.

7. Dirac operators and ordinary finite spaces

Connes has shown [C1, C2, C3] that for a compact Riemannian (spin) manifold all the metric information is contained in the Dirac operator. This led him to suggest that for “non-commutative spaces”, metrics should be specified by some analogue of Dirac operators. We explore here some aspects of this suggestion for finite-dimensional commutative C^{∗}-algebras, i.e. ordinary finite spaces. This will clarify some of the considerations of the previous sections. Here and throughout all the rest of this paper, when we say that an operator D is a “ Dirac” operator, this is not meant to indicate any particular properties of D, but rather is meant to indicate how D is employed, namely to define a Lipschitz seminorm. Let X be a finite set, and let A = C(X). In order to remain fully in the setting of the previous sections we take C(X) to consist only of real-valued functions. But in the present commutative situation this is not so important because, unlike the non-commutative case, if one does not know the algebra structure, the norm for complex-valued functions is still given by a simple formula in terms of the norm for real-valued functions. (See e.g. lemma 14 of [W2].) Consequently we will be a bit careless here about this distinction. We will suppose that A has been faithfully represented on a finite-dimensional complex Hilbert space H. We suppose given on H an operator D (the “Dirac” operator). It is usual to take D to be self-adjoint. But we find it slightly more convenient to take D to be skew-adjoint. The two choices are related by a multiplication by i, and give the same metric results. Following Connes, we define a seminorm, L, on A by

L(a) = ∥[D, a]∥,

where [ , ] denotes the usual commutator of operators, and the norm is the operator norm. We want L to be a Lip-norm. Thus we require that if [D, a] = 0 then a ∈ ℂI. Because we are in a finite-dimensional setting, L is continuous for ∥ ∥_{∞}, and indeed is a Lip-norm on A. From L we obtain a metric, ρ_{L}, on the space S(A) of probability measures on X, as well as on its set of extreme points, which is identified with X itself. We now give a very simple example to show that ρ_{L} on S(A) need not agree with the metric obtained from ρ_{L} on X.

7.1 Example. Consider a three-dimensional commutative C^{∗}-algebra, A, represented faithfully on a three-dimensional Hilbert space. Thus we can identify A with the algebra of diagonal matrices in the full matrix algebra M_{3} = M_{3}(ℂ). We will consider Dirac operators of a special form which facilitates calculation, namely matrices D in M_{3}(ℂ) of the form

D = ( \begin{matrix}0 & 0 & α \\ 0 & 0 & β \\ −α & −β & 0 \end{matrix} )

where α > 0 and β > 0. We will also restrict to those f ∈ A which are real, and denote the three values (or diagonal entries) of f by (f_{1}, f_{2}, f_{3}). Because D is skew-symmetric, [D, f] is a real symmetric matrix, whose eigenvalues thus are real. In fact, we have

[D, f] = ( \begin{matrix}0 & 0 & α(f_{3} − f_{1}) \\ 0 & 0 & β(f_{3} − f_{2}) \\ α(f_{3} − f_{1}) & β(f_{3} − f_{2}) & 0 \end{matrix} ) .

Because of this special form, the eigenvalues are easily calculated, and one finds that

L(f) = ∥[D, f]∥ = (α^{2}(f_{3} − f_{1})^{2} + β^{2}(f_{3} − f_{2})^{2})^{1/2}.

It is clear from this that if L(f) = 0 then f is a constant function. Thus L defines a Lip-norm on A. We now proceed to calculate the corresponding metric on S(A). We first calculate the dual norm, L′, on A′^{0}, the dual space of Ã, with notation as in the previous sections. We identify A′^{0} with real diagonal matrices of trace 0, paired with A via the trace. For λ ∈ A′^{0} we denote its components by λ = (λ_{1}, λ_{2}, λ_{3}). Of course

L′(λ) = sup{|⟨f, λ⟩| : L(f) ≤ 1}.

Now both |⟨f, λ⟩| and L(f) are unchanged if we add a constant function to f. Thus for the supremum defining L′(λ) we can assume that f_{3} = 0 always.

Furthermore, we know that λ_{3} = −(λ_{1} + λ_{2}). Thus we need only deal with the first two components of f and λ. We do this without changing notation. Then we see that

L′(λ) = sup{|f_{1}λ_{1} + f_{2}λ_{2}| : α^{2}f_{1}^{2} + β^{2}f_{2}^{2} ≤ 1}.

But this is just the norm of a functional on a suitable Hilbert space. Specifically, let l^{2}(w) be the Hilbert space of functions on a 2-point space with weight function w given by (α^{2}, β^{2}). Then

f_{1}λ_{1} + f_{2}λ_{2} = f_{1}(λ_{1}/α^{2})α^{2} + f_{2}(λ_{2}/β^{2})β^{2},

and in this form the norm of the functional is the length of the vector in l^{2}(w) defining it. This gives

L′(λ) = ((λ_{1}/α^{2})^{2}α^{2} + (λ_{2}/β^{2})^{2}β^{2})^{1/2}
= (λ_{1}^{2}/α^{2} + λ_{2}^{2}/β^{2})^{1/2}.

We now apply this to obtain the metric on S(A). If μ, ν ∈ S(A), then for the evident notation

ρ_{L}(μ, ν) = L′(μ − ν) = ((μ_{1} − ν_{1})^{2}/α^{2} + (μ_{2} − ν_{2})^{2}/β^{2})^{1/2}.

Let X denote the maximal ideal space of A. We identify its 3 points with the 3 extreme points of S(A), and label them, corresponding to the coordinates in A, by δ_{1}, δ_{2}, δ_{3}. Then from the above formula for ρ_{L} we find that the metric on X is given by:

ρ_{L}(δ_{1}, δ_{2}) = (1/α^{2} + 1/β^{2})^{1/2}
ρ_{L}(δ_{1}, δ_{3}) = 1/α
ρ_{L}(δ_{2}, δ_{3}) = 1/β.

Define γ by ρ_{L}(δ_{1}, δ_{2}) = 1/γ. Let L^{e} denote the ordinary Lipschitz norm on A coming from this metric on X. Then

L^{e}(f) = max{|f_{1} − f_{2}|γ, |f_{1} − f_{3}|α, |f_{2} − f_{3}|β}.

Clearly L^{e} is quite different from L. From Theorem 4.1 we know that the metrics on S(A) will thus be quite different, even though they agree on the extreme points. This is, of course, also easily seen by direct calculations.

We now make some observations in preparation for the next section. It is wellknown [W1, W2] that the Lipschitz seminorms L = L_{ρ} from ordinary metrics on a metric space X have a nice relation to the lattice structure of (real-valued) C(X), namely

L(f ∨ g) ≤ L(f) ∨ L(g).

We remark that for the L of the above example this inequality fails. For instance, with notation as above, let f = (1, 0, 0) and g = (0, 1, 0), so that f ∨ g = (1, 1, 0). Then we see that

L(f) = α, L(g) = β, while L(f ∨ g) = (α^{2} + β^{2})^{1/2}.

(This is related to the counterexample following theorem 16 of [W2].) However, it is not difficult to check that the above L does satisfy the weaker inequality

L(f ∨ 0) ≤ L(f).

In fact, one can prove that this holds for any choice of skew-adjoint D for the above A. To find a counterexample for this weaker inequality one must take A to be 4-dimensional. I have not found a systematic way of constructing a counterexample there, but some examination of what is needed, followed by some experimentation with MATLAB yields the following (and related) example:

D = ( \begin{matrix}0 & 4 & −1 & 0 \\ −4 & 0 & 2 & −2 \\ 1 & −2 & 0 & −4 \\ 0 & 2 & 4 & 0 \end{matrix} )

and f = (4, 2, 0, −1). We remark that ordinary Lipschitz norms on compact metric spaces can all be easily obtained by means of Dirac operators. I pointed this out in a lecture in 1993, and the details are indicated after the proof of proposition 8 of [W2]. See also the discussion for graphs which we will give toward the end of Section 11.

8. A characterization of ordinary Lipschitz seminorms

Let X be a compact space, let ρ be a metric on X (giving the topology of X), and let L denote the corresponding ordinary Lip-norm on C(X) (permitted to take value +∞). As just mentioned in the last section, it is well-known [W1, W2] and easy to prove that L relates nicely to the lattice structure of C(X) by means of the inequality

L(f ∨ g) ≤ L(f) ∨ L(g).

In Weaver’s more general setting of domains of W^{∗}-derivations he proves this inequality for W^{∗}-derivations of Abelian structure. (See lemma 12 of [W2].) We show here that the above inequality exactly characterizes the Lip-norms which are the ordinary Lipschitz seminorms coming from ordinary metrics on X. We remark that we never assume here that our Lip-norms satisfy the Leibniz inequality for the algebra structure, namely

L(fg) ≤ L(f)∥g∥ + ∥f∥L(g).

But ordinary Lipschitz seminorms do satisfy this inequality. Thus one consequence of this section is that the above lattice inequality implies the Leibniz inequality. On the other hand, the Lip-norm from any “Dirac” operator will satisfy the Leibniz inequality, but can easily fail to satisfy the lattice inequality, as we saw by examples in the previous section. Thus the lattice inequality is much stronger than the Leibniz inequality. However we should point out that for Dirac operators on compact spin Riemannian manifolds, in spite of their being defined by means of various partial derivatives and spinors, the corresponding Lip-norms do satisfy the lattice inequality. This is because Connes shows [C1, C2, C3] that the Lip-norms which those Dirac operators define coincide with the ordinary Lip-norms for the ordinary metrics on the manifolds determined by the Riemannian metrics. Recall that for us C(X) consists of real-valued functions.

8.1 Theorem. Let X be a compact space, let A be a dense subspace of C(X) containing the constant functions, and let L be a Lip-norm on A. Let L̄ denote the closure of L, viewed as defined on all of C(X) as in the discussion before Proposition 4.4, and thus permitted to take value +∞. Then the following conditions are equivalent:

1. The Lip-norm L is the restriction to A of the usual Lipschitz seminorm corresponding to a metric on X (namely the metric ρ_{L}).

2. For every f, g ∈ C(X) we have

L̄(f ∨ g) ≤ L̄(f) ∨ L̄(g).

The following lemma is somewhat parallel to lemma 13 of [W2]. For later use we state it in slightly greater generality than needed immediately.

8.2 Lemma. Let A be a dense subspace of C(X) containing the constant functions, and closed under the finite lattice operations (i.e. if f, g ∈ A then f ∨ g ∈ A). Let L be a Lip-norm on A which satisfies the inequality

L(f ∨ g) ≤ L(f) ∨ L(g)

for all f, g ∈ A. Let L̄ be the closure of L, defined on all of C(X), permitted to take value +∞. Let F be a bounded subset of A for which there is a constant, k, such that L(f) ≤ k for all f ∈ F. Let g = sup{f ∈ F}. Then g ∈ C(X) and L̄(g) ≤ k.

Proof. Let {g_{α}} be the net of suprema of finite subsets of F. Then {g_{α}} is contained in A, and converges up to f pointwise. By the hypothesis on L we have L(g_{α}) ≤ k for all α. Thus we have

|g_{α}(x) − g_{α}(y)| ≤ kρ_{L}(x, y)

for all α and all x, y ∈ X; that is, {g_{α}} is equicontinuous. We can thus apply the Ascoli theorem [Ru] to conclude that the net {g_{α}} has a subnet which converges uniformly. But the limit of this subnet must be g, and so g must be continuous. Furthermore, from the lower semicontinuity of L̄ we must have L̄(g) ≤ k. □

Proof of Theorem 8.1. As indicated above, it is basically well-known, and not hard to verify, that condition 1 implies condition 2. Suppose conversely that condition 2 holds. For any x ∈ X let ρ_{L}^{x} be the continuous function on X defined by ρ_{L}^{x}(y) = ρ_{L}(x, y). Set S_{x} = {f ∈ A : f(x) = 0, L(f) ≤ 1}. Since L(f) is unchanged when a constant function is added to f, or when f is replaced by −f, the definition of ρ_{L} can be rewritten as

ρ_{L}^{x}(y) = sup{f(y) : f ∈ S_{x}}.

This means that ρ_{L}^{x} = sup S_{x}. But S_{x} is a bounded set in A by the finite radius considerations. Thus we can apply the above lemma to conclude that L̄(ρ_{L}^{x}) ≤ 1. Suppose that L̄(ρ_{L}^{x}) = c < 1. Then L̄((1/c)ρ_{L}^{x}) = 1, and so from the definition of ρ_{L} we obtain

(1/c)|ρ_{L}^{x}(x) − ρ_{L}^{x}(y)| ≤ ρ_{L}(x, y),

for all y ∈ X, that is,

ρ_{L}(x, y) ≤ cρ_{L}(x, y)

for all y ∈ X, which is impossible (unless X has only one point, which we now do not permit). Thus L̄(ρ_{L}^{x}) = 1. Much as in Section 6, let L^{e} denote the ordinary Lip-norm on C(X) (permitting value +∞) corresponding to the restriction of ρ_{L} as metric on X. (Recall that X is identified with the extreme points of S(A).) As seen in Proposition 6.1, L^{e} ≤ L̄. We now show that L^{e} = L̄ because of the inequality in the hypotheses of our theorem (and its extension in Lemma 8.2). Let f ∈ C(X), and suppose that L^{e}(f) ≤ 1. Thus

|f(x) − f(y)| ≤ ρ_{L}(x, y)

for all x, y ∈ X. In particular

f(x) − ρ_{L}(x, y) ≤ f(y).

For each x ∈ X define h^{x} ∈ C(X) by

h^{x}(y) = f(x) − ρ_{L}(x, y).

Then the above inequality says that h^{x} ≤ f for each x. But it is clear that h^{x}(x) = f(x). Thus f = sup{h^{x} : x ∈ X}. Then from the considerations of the previous paragraph we see that L̄(h^{x}) = 1 for all x. Thus by Lemma 8.2 we have L̄(f) ≤ 1. It follows that L̄ = L^{e} as desired. □

8.3 Corollary. Let X be a compact space, and let A be a dense subspace of C(X) which contains the constant functions and is closed under the finite lattice operations. Let L be a Lip-norm on A, and suppose that

L(f ∨ g) ≤ L(f) ∨ L(g)

for all f, g ∈ A. Then L is the restriction to A of the ordinary Lip-norm on C(X) corresponding to the metric ρ_{L} on X.

Proof. Let f, g ∈ C(X). Then from Lemma 8.2 we see immediately that

L̄(f ∨ g) ≤ L̄(f) ∨ L̄(g).

We can thus apply Theorem 8.1 to obtain the desired conclusion. □

One way of viewing Theorem 8.1 is that it characterizes the Lip-norms on commutative C^{∗}-algebras which come from the corresponding metric on the extreme points of S(A). It would be interesting to have a corresponding characterization for non-commutative C^{∗}-algebras, and for general order-unit spaces.

9. Lip-norms from metrics on S(A)

It is natural to ask which metrics on S(A) arise from Lip-norms on A. We obtain here a characterization of such metrics. Many of the steps work for arbitrary convex sets, and so at first we will work in that setting. Thus we let V be any vector space over ℝ, and we let K be any convex set in V which spans V. Much as above, let D_{2} = K − K. Note that not only is D_{2} convex, but it is also balanced, in the sense that if λ ∈ D_{2} and if t ∈ [−1, 1], then tλ ∈ D_{2}. To see this, note that if λ ∈ D_{2} then clearly −λ ∈ D_{2}, so we only need consider t ≥ 0. But

t(μ − ν) = μ − (tν + (1 − t)μ),

which is in D_{2} by the convexity of K. Let V^{0} = ℝD_{2}. Then V^{0} is a vector subspace of V. In the setting where K = S(A) we know that V^{0} is a proper subspace of V. Let M be a norm on V^{0}. Then we can define a metric, ρ, on K by ρ(μ, ν) = M(μ − ν). We want to characterize the metrics which arise in this way. The most natural property to expect is that ρ be convex (in each variable), that is:

9.1 Definition. We say that a metric ρ on K is convex if for every μ, ν_{1}, ν_{2} ∈ K and t ∈ [0, 1] we have

ρ(μ, tν_{1} + (1 − t)ν_{2}) ≤ tρ(μ, ν_{1}) + (1 − t)ρ(μ, ν_{2}).

The metrics coming from norms on V^{0} are convex because

μ − (tν_{1} + (1 − t)ν_{2}) = t(μ − ν_{1}) + (1 − t)(μ − ν_{2}).

Given a metric ρ on K, our strategy will be to try to use ρ to define a norm, M, on V^{0} by first defining it on D_{2}. Specifically, for λ ∈ D_{2} we would like to set

M(λ) = ρ(μ, ν)

for λ = μ − ν with μ, ν ∈ K. But we need to know that this is well-defined. That is, we need to know that if μ, ν, μ′, ν′ ∈ K and if μ − ν = μ′ − ν′, then ρ(μ, ν) = ρ(μ′, ν′). This can be rewritten in terms of midpoints so as to appear a bit closer to considerations of convexity, namely, that if

(9.2)  (μ + ν′)/2 = (μ′ + ν)/2

then ρ(μ, ν) = ρ(μ′, ν′). This clearly holds for the metrics coming from norms. One finds an attractive geometrical interpretation when one draws a picture of this relation.

9.3 Definition. We say that a metric ρ on K is midpoint-balanced if whenever equation (9.2) above holds, it follows that ρ(μ, ν) = ρ(μ′, ν′).

Let us now assume that ρ is midpoint-balanced. Then M on D_{2} is well-defined as above. We wish to extend it to a norm on V^{0}. For this to be possible we first must have the property that if t ∈ ℝ, |t| ≤ 1, and if λ ∈ D_{2}, then M(tλ) = |t|M(λ). Now from the definition of M it is clear that M(−λ) = M(λ). Thus it suffices to treat the case in which t ≥ 0. If λ = μ − ν, then

tλ = t(μ − ν) = μ − (tν + (1 − t)μ),

so that by the definition of M we have M(tλ) = ρ(μ, tν + (1 − t)μ). From convexity, ρ(μ, tν +(1−t)μ) ≤ tρ(μ, ν). But also tλ = (tμ+(1−t)ν)−ν, which gives a similar inequality. Then from the triangle inequality and convexity we have

ρ(μ, ν) ≤ ρ(μ, tν + (1 − t)μ) + ρ(tν + (1 − t)μ, ν)
≤ tρ(μ, ν) + (1 − t)ρ(μ, ν) = ρ(μ, ν).

Thus the inequalities must be equalities, and we obtain:

9.4 Lemma. Let ρ be a metric on K which is convex and midpoint balanced. Define M on D_{2} as above using ρ. Then for any μ, ν ∈ S(A) and t ∈ [0, 1] we have

ρ(μ, tν + (1 − t)μ) = tρ(μ, ν),

and for any λ ∈ D_{2} and t ∈ [−1, 1] we have

M(tλ) = |t|M(λ).

Next, we need that M is subadditive on D_{2}. This means that if λ, λ′ ∈ D_{2} and if λ+λ′ ∈ D_{2}, then M(λ+λ′) ≤ M(λ)+M(λ). Let λ = μ−ν, λ′ = μ′−ν′. Then λ + λ′ = (μ + μ′) − (ν + ν′). Assuming that ρ is convex and midpoint-balanced, we obtain from Lemma 9.4 that

M(λ + λ′) = 2M((λ + λ′)/2).

Now (λ + λ′)/2 = (μ + μ′)/2 − (ν + ν′)/2, and (μ + μ′)/2, (ν + ν′)/2 ∈ S(A). Thus

M((λ + λ′)/2) = ρ((μ + μ′)/2, (ν + ν′)/2),

and we see that what we need is:

9.5 Definition. We say that a metric ρ on K is midpoint concave if for any μ, ν, μ′, ν′ ∈ K we have

ρ((μ + μ′)/2, (ν + ν′)/2) ≤ (1/2)(ρ(μ, ν) + ρ(μ′, ν′)).

Again one finds an attractive geometrical interpretation when one draws a picture of this inequality. From the discussion above we now know that:

9.6 Lemma. Let ρ be a metric on K which is convex, midpoint balanced, and midpoint concave. Define M on K as above. If λ, λ′ ∈ D_{2} and if λ + λ′ ∈ D_{2}, then

M(λ + λ′) ≤ M(λ) + M(λ′).

9.7 Theorem. Let ρ be a metric on the convex subset K of V, and let V^{0} = ℝD_{2} = ℝ(K − K). Then there is a norm, M, on V^{0} such that ρ(μ, ν) = M(μ − ν) for all μ, ν ∈ K, if and only if ρ is convex, midpoint balanced, and midpoint concave. The norm M is unique.

Proof. The uniqueness is clear since V^{0} = ℝ(K −K). We have seen above that the conditions on ρ are necessary. We now show that they are sufficient. We let M be defined on D_{2} = K − K as above. For any λ ∈ V^{0} there is a t > 0 such that tλ ∈ D_{2}. We want to extend M to V^{0} by setting

M(λ) = t^{−1}M(tλ).

From Lemma 9.4 it is easily seen that M is well-defined, and furthermore that M(sλ) = |s|M(λ) for all s ∈ ℝ and λ ∈ V^{0}. The subadditivity of M then follows easily from Lemma 9.6. □

We now want to apply the above ideas to S(A) for an order-unit space A. Note that the V^{0} of just above is then the A′^{0} of earlier. We will need the following theorem, which does not involve the above ideas.

9.8 Theorem. Let A be an order-unit space, and let M be a norm on A′^{0}. Define a metric, ρ, on S(A) by

ρ(μ, ν) = M(μ − ν).

If the ρ-topology coincides with the weak-∗ topology on S(A), then

M = (L_{ρ})′

on A′^{0}.

Proof. Since Lip_{ρ} is a subspace of C(S(A)), we can set A_{L} = (Lip_{ρ})∩Af(S(A)). Note that A_{L} need not be contained in A unless A is complete. Initially it is not clear how big A_{L} is. Parallel to our earlier notation, let V denote the normed space A′^{0} with norm M. Note that V need not be complete. Let V′ denote the Banach space dual of V, with dual norm M′. Fix any ν_{0} ∈ S(A). For any φ ∈ V′ define a function, τ(φ), on S(A) by

τ(φ)(μ) = φ(μ − ν_{0}).

Then for μ, ν ∈ S(A) we have

|τ(φ)(μ) − τ(φ)(ν)| = |φ(μ − ν)| ≤ M′(φ)M(μ − ν) = M′(φ)ρ(μ, ν).

Thus τ(φ) ∈ Lip_{ρ} and L_{ρ}(τ(φ)) ≤ M′(φ). In particular, τ(φ) is continuous on S(A) since ρ gives the weak-∗ topology. Furthermore it is easily seen that τ(φ) is affine on S(A). Thus τ(φ) ∈ A_{L}. Consequently τ is a norm-non-increasing linear map from (V′, M′) to (A_{L}, L_{ρ}). Let τ̃ denote τ composed with the map from A_{L} to Ã_{L}. Then it is easily seen that τ̃ does not depend on the choice of ν_{0}. We now need:

9.9 Lemma. Let Ā = Af(S(A)), the completion of A for ∥ ∥, so that A_{L} ⊆ Ā. Then A_{L} is dense in Ā.

Proof. Since ℝe ⊆ A_{L}, it suffices to show that Ã_{L} is dense in Ā^{∼}. Let λ ∈ D_{2} ⊆ A′^{0} = (Ā^{∼})′. Suppose that λ(A_{L}) = 0. Let λ = μ − ν with μ, ν ∈ S(A). For any φ ∈ V′ we have τ(φ) ∈ A_{L}, so

0 = λ(τ(φ)) = μ(τ(φ)) − ν(τ(φ)) = φ(μ − ν_{0}) − φ(ν − ν_{0}) = φ(λ).

Since this is true for all φ ∈ V′, it follows that λ = 0. Since D_{2} spans A′^{0}, an application of the Hahn–Banach theorem now shows that A_{L} is dense on Ā. □

Now let f ∈ A_{L}. We seek to define a linear functional, σ(f), on A′^{0} related to the σ in the proof of Theorem 5.2. We first try to define σ on D_{2} by

σ(f)(λ) = f(μ) − f(ν),

where λ = μ − ν for μ, ν ∈ S(A). But we need to show that σ(f) is welldefined. We argue much as we did before Definition 9.3. If also λ = μ_{1} − ν_{1} for μ_{1}, ν_{1} ∈ S(A), then (μ + ν_{1})/2 = (μ_{1} + ν)/2. But these are elements of S(A) and so

f((μ + ν_{1})/2) = f((μ_{1} + ν)/2).

But from the fact that f is affine it now follows that

f(μ) − f(ν) = f(μ_{1}) − f(ν_{1}).

Thus σ(f) is well-defined on D_{2}. We now need to know that σ(f) is “linear” on D_{2}. The proof that σ(f)(tλ) = tσ(f)(λ) for t ∈ [−1, 1] is similar to the proof of Lemma 9.4. The proof that σ(f)(λ + λ_{1}) = σ(f)(λ) + σ(f)(λ_{1}) if λ + λ_{1} ∈ D_{2} is similar to the argument just before Definition 9.5. The proof that σ(f) then extends to a linear functional on A′^{0} is similar to the arguments in the proof of Theorem 9.7. For λ = μ − ν with μ, ν ∈ S(A) we have

|σ(f)(λ)| = |f(μ) − f(ν)| ≤ L_{ρ}(f)ρ(μ, ν) = L_{ρ}(f)M(μ − ν) = L_{ρ}(f)M(λ).

It follows that σ(f) ∈ V′ and M′(σ(f)) ≤ L_{ρ}(f). Thus σ is a norm-non-increasing linear map from (A_{L}, L_{ρ}) to (V′, M′). Note that the constant functions are in the kernel of σ, so that σ determines a norm-non-increasing linear map from (Ã_{L}, L̃_{ρ}) to (V′, M′). But for f ∈ A_{L} we have

τ(σ(f))(μ) = σ(f)(μ − ν_{0}) = f(μ) − f(μ_{0}).

Consequently τ̃(σ̃(f̃)) = f̃. Similarly, for φ ∈ V′ and λ = μ − ν we have

σ̃(τ̃(φ))(λ) = τ(φ)(μ) − τ(φ)(ν) = φ(μ − ν_{0}) − φ(ν − ν_{0}) = φ(λ),

so that σ̃(τ̃(φ)) = φ. Thus σ̃ and τ̃ are inverses of each other. Since they are norm-non-increasing, we obtain:

9.10 Lemma. The map τ̃ is an isometric isomorphism of (V′, M′) onto (A_{L}, L_{ρ}), with inverse σ̃.

We can now complete the proof of Theorem 9.8. Since A_{L} is dense in Ā by Lemma 9.9, for any λ ∈ V′ we have

(L_{ρ})′(λ) = sup{λ(τ̃(φ)) : L_{ρ}(τ̃(φ)) ≤ 1} = sup{φ(λ) : M′(φ) ≤ 1} = M(λ).

□

Putting together the various pieces of this section, we obtain:

9.11 Theorem. Let A be an order-unit space, and let ρ be a metric on S(A) which gives the weak-∗ topology. Then ρ comes from a Lip-norm L on A via the relation

ρ(μ, ν) = L′(μ − ν)

if and only if ρ is convex, midpoint balanced, and midpoint convex.

Nik Weaver has suggested to me the following alternative treatment of the material of this section. Let V, K, and V^{0} be as at the beginning of this section.

9.12 Definition. We say that a metric ρ on K is linear if for every μ, ν ∈ K, every v ∈ V^{0}, and every t ∈ ℝ^{+} such that μ + tv and ν + v are in K we have

ρ(μ, μ + tv) = tρ(ν, ν + v).

It is easily seen that if ρ comes from a norm on V^{0} then ρ is linear. Conversely, if ρ is linear, define a norm, M, on V^{0} by

M(v) = ρ(μ, μ + tv)/t

for any μ ∈ K and any t ∈ ℝ^{+} such that μ + tv ∈ K. One checks that M is well-defined and is indeed a norm. Furthermore, ρ comes from M. Weaver also points out that if V is a locally convex topological vector space and if K is compact, then for a suitable definition of ρ being compatible with the topology, one can show that when ρ is linear and compatible, then K is isometrically isomorphic to S(Af(K)) when the latter is given the metric coming from the Lipschitz seminorm on Af(K) coming from ρ. It is not clear that examples will come up where it is actually useful to apply the considerations of this section in order to obtain Lip-norms. Until such examples arise, it will not be clear whether my version or Weaver’s will be the more useful.

10. Musings on metrics

Since the theory in the previous sections worked for order-unit spaces, which need not be algebras, the Leibniz inequality played no significant role there. Indeed, even when one has an algebra, I have not seen how to make effective use of the Leibniz inequality. Nevertheless, most constructions of Lipschitz seminorms which I have seen in the literature seem to provide ones which do satisfy the Leibniz inequality. We will briefly explore here a variety of such constructions, and the relationships between them. Our interest will be on seeing general patterns, and we will not try to deal carefully with the many technical issues which arise. Thus we will be less precise than in the previous sections. A very natural way to look for Lipschitz seminorms, closely related to Weaver’s W^{∗}-derivations [W2], goes as follows. Let A be a unital algebra and let (Ω, d) be a first-order differential calculus for A. Thus Ω (which is also often denoted Ω^{1}) is an A-A-bimodule, and d is an Ω-valued derivation on A, that is, a linear map from A into Ω which satisfies the Leibniz identity

d(ab) = (da)b + a(db).

We do not require that the range of d generates Ω. Suppose now that A is in fact a normed algebra, and that we have a bimodule norm, N, on Ω (for the norm ∥ ∥ on A), that is, a norm such that

N(aωb) ≤ ∥a∥N(ω)∥b∥

for a, b ∈ A and ω ∈ Ω. Define a seminorm L on Ω by

L(a) = N(da).

It is easily seen that L satisfies the Leibniz inequality. Since d1 = 0, we have L(1) = 0. Of course, without further hypotheses the null-space of L may be much bigger. (We should mention that not all seminorms satisfying the Leibniz inequality can be constructed in this way—see the discussion in [BC].) There is a universal first-order differential calculus for any unital algebra A [Ar, C2]. We approach this in a way which emphasizes more than usual those differential calculi which are inner, since at least conceptually that is what Dirac operators give, as we will see shortly. We form the algebraic tensor product

Ω_{1}^{u} = A ⊗ A,

with bimodule structure defined as usual by a(b ⊗ c)d = ab ⊗ cd. We define d by

da = 1 ⊗ a − a ⊗ 1.

10.1 Definition. A first-order calculus (Ω, d) is inner if there is a ω_{0} ∈ Ω such that

da = ω_{0}a − aω_{0}.

Then the calculus (Ω_{1}^{u}, d) defined above is inner, with ω_{0} = 1 ⊗ 1. Note that here ω_{0} may not be in the sub-bimodule generated by the range of d. This is an indication of why we do not require this generation property. It is simple to verify:

10.2 Proposition. The inner first-order calculus (Ω_{1}^{u}, d, 1 ⊗ 1) is universal among inner first-order differential calculi over A, in the sense that if (Ω′, d′, ω′_{0}) is any other inner first-order differential calculus, then there is a bimodule homomorphism Φ : Ω_{1}^{u} → Ω′ such that Φ(da) = d′a and Φ(1 ⊗ 1) = ω′_{0}. In particular,

Φ(a ⊗ b) = aω′_{0}b

for a, b ∈ A. If Ω′ is generated by ω′_{0} as bimodule, then Φ is surjective, so that Ω′ is a quotient of Ω_{1}^{u}.

10.3 Proposition. Any first-order differential calculus is contained in an inner first-order calculus.

Proof. Let (Ω, d) be a first-order calculus. Set Ω̄ = Ω ⊕ A as left A-module, set d̄a = da ⊕ 0, and set ω̄_{0} = 0 ⊕ 1. We must extend the right action of A on Ω to a right action on Ω̄ such that d̄a = ω̄_{0}a − aω̄_{0}. Thus it is clear that we must set (0 ⊕ 1)a = ω̄_{0}a = da ⊕ 0 + aω̄_{0} = da ⊕ a, and so

(ω, b)a = (ωa + bda, ba).

It is simple to check that this gives the desired structure. □

Now let Ω^{u} denote the sub-bimodule of Ω_{1}^{u} generated by the range of d, and so spanned by elements of the form

adb = a ⊗ b − ab ⊗ 1.

Let (Ω′, d′) be a first-order differential calculus which is not inner. Expand it to an inner calculus by the construction of the previous proposition, and then restrict Φ of that proposition to Ω^{u}. It is clear from the construction that Φ will carry Ω^{u} into Ω′, where Ω′ is viewed as a sub-bimodule of its expansion. We obtain in this way:

10.4 Proposition. The calculus (Ω^{u}, d) is universal among all first-order differential calculi over A, in the sense that if (Ω′, d′) is any other first-order differential calculus, then there is a bimodule homomorphism Φ : Ω^{u} → Ω such that Φ(da) = d′a. If Ω′ is generated by the range of d′ as bimodule, then Φ is surjective, so that Ω′ is a quotient of Ω^{u}.

We notice that if (Ω, d) is any first-order differential calculus and if N is any subbimodule of Ω, then we obtain a calculus (Ω/N , d′) where d′ is the composition of d with the canonical projection of Ω onto Ω/N. However, unlike the universal calculus, there may now be many more elements a for which da = 0 beyond the scalar multiples of 1. Let us examine briefly what the above looks like when A = C(X) for a compact space X. Then Ω_{1}^{u}(= A ⊗ A) is naturally viewed as a dense sub-bimodule, in fact subalgebra, of C(X × X). The bimodule actions are, of course,

(fF)(x, y) = f(x)F (x, y), (F f)(x, y) = F (x, y)f(y),

and ω_{0} = 1 ⊗ 1 is the constant function 1, so that d is given by

(df)(x, y) = f(y) − f(x).

Then Ω^{u} is spanned by the fdg, where

(fdg)(x, y) = f(x)(g(y) − g(x)).

Thus the elements of Ω^{u} take value 0 on the diagonal, Δ, of X × X, and consequently Ω^{u} ⊆ C_{∞}(X × X ∖ Δ). In fact it is easy to see that Ω^{u} is a dense subalgebra of C_{∞}(X × X ∖ Δ). Let ρ be an ordinary metric on X (giving the topology of X). View ρ as a strictly positive function on X × X ∖ Δ, and let γ = ρ^{−1}. Then γ is a continuous function on X × X ∖ Δ, but γ is unbounded if X is not finite. Let C(X × X ∖ Δ) denote the algebra of continuous possibly-unbounded functions on X × X ∖ Δ. Then C(X × X ∖ Δ) can be viewed as the algebra of operators affiliated with the C^{∗}-algebra C_{∞}(X ×X ∖Δ) in the sense studied by Baaj [Ba] and Woronowicz [Wo]. In an evident way C(X × X ∖ Δ) is an A-A-bimodule, containing γ.

There are now two routes which we can take. One is to consider the innerderivation, d_{γ}, defined by γ. Thus

(d_{γ}f)(x, y) = γ(x, y)f(y) − f(x)γ(x, y) = (f(y) − f(x))/ρ(x, y).

Then we can consider bimodule norms, possibly taking value +∞, on C(X × X ∖ Δ), as a way to obtain Lipschitz norms on A. The other route is to use γ (or ρ) to directly define norms on C_{∞}(X × X ∖ Δ). For the first route the most obvious norm is the supremum norm, which leads to the usual definition of the Lipschitz seminorm for a metric space. However, we choose to explore further the second route. (But most of what we find will have a fairly evident reinterpretation in terms of the first route.) There is a large variety of ways to obtain bimodule norms on C_{∞}(X × X ∖ Δ). The one which gives the usual definition of the Lipschitz seminorm for a metric is clearly

N(F) = ∥γF ∥_{∞},

permitted to take value +∞. But here are some others. Let m be any positive (finite) measure on X, and assume that m × m restricted to X × X ∖ Δ has as support all of X × X ∖ Δ. Then one can consider all of the L^{p}-norms for m×m. If one wants to put γ (or ρ) explicitly into the picture, one can consider the measure γ(m × m), although this just represents the choice of a different measure. Note that if f is an ordinary Lipschitz function for ρ, then γdf is a bounded function on X × X ∖ Δ, so that ∥γdf∥_{p,m×m} is finite. Thus the subalgebra of elements of A for which this Lipschitz seminorm is finite is dense in A. To explore further possibilities, let us for simplicity assume that X is finite. Then Ω_{1}^{u} = C(X ×X) can be viewed as the algebra of all matrices whose entries are indexed by elements of X × X. The left and right actions of A on Ω_{1}^{u} can be viewed as coming from embedding A as the diagonal matrices and using left and right matrix multiplication. Then ω_{0} is the matrix with a 1 in each entry. On A we keep the supremum norm, but on the matrix algebra Ω_{1}^{u} we can consider any A-A-bimodule norm. Let B denote Ω_{1}^{u} viewed as matrix algebra, and equipped with the usual C^{∗}-algebra norm. View Ω_{1}^{u} as a B-B-bimodule in the evident way. Then we can consider B-B-bimodule norms on Ω_{1}^{u}. Any such will in particular be an A-A-bimodule norm. But there has been extensive study of the possible B-B-bimodule norms on Ω_{1}^{u}. They are commonly called “symmetric norms”, and among the best known are the Schatten p-norms, which include the Hilbert–Schmidt norm and the trace norm. These have, of course, also been extensively studied for operators on infinite dimensional Hilbert spaces, and play a fundamental role in Connes’ theory of integration on non-commutative spaces. (See [C2] Chapter IV and its Appendix D. A nice treatment of the finite case can be found in [Bh].) From every symmetric norm we obtain a Lip-norm on A (since A is finite-dimensional). This does not exhaust the possibilities, as there is no necessity to restrict to symmetric norms in order to get A-A-bimodule norms.

All of the above discussion has been for the universal differential calculus. We get many more possibilities by using other differential calculi. We continue to concentrate on the case of A = C(X) with X compact. Now sub-A-A-bimodules of C(X × X), when closed in the supremum norm, will be ideals of C(X ×X), and the quotient can be identified with C(W) for some closed subset W of X × X. We can restrict df to W. But some condition must be placed on W if we want to ensure that df|_{W} = 0 only if f is a constant function. For this purpose it is convenient to assume, to begin with, that W contains the diagonal Δ and is symmetric about Δ, that is, if (x, y) ∈ W then (y, x) ∈ W. Given x ∈ X we define the W -neighborhood of x to be the (closed) set of those y ∈ X such that (x, y) ∈ W. By the W -component of x we mean the smallest closed subset of X which contains the W -neighborhood of each of its points. If df|_{W} = 0, then f is constant on the W -component of each point. Thus a sufficient condition under which df|_{W} = 0 will imply that f is constant, is that the W -component of each point is all of X. If X is a finite set, then W ∖Δ can be viewed as consisting of the directed edges for a graph whose vertices are the points of X. Then the above condition becomes the condition that this graph is connected in the usual sense. If X is not discrete, it is usual to require that W is a neighborhood of Δ. Then each W -neighborhood of a point will be an ordinary (closed) neighborhood, and so the W -component of each point will be both closed and open. In particular, if X is connected it will be true that df|_{W} = 0 implies that f is constant. We remark that if W is a neighborhood of Δ and is symmetric about Δ, and if we set Ω = C(W), then the first order calculus (Ω, d) obtained as above is the typical degree-one piece of the complexes (Ω_{W}^{∗} , d) used in defining the Alexander–Spanier cohomology of X. The higher-degree pieces are defined similarly but in terms of X^{n} for various n. The Alexander–Spanier cohomology is then obtained by taking a limit of the homology of these complexes as W shrinks to Δ. Essentially this view can be seen in lemma 1.1 of [CM], where smooth functions on a manifold are used, and in Section 1 of [MW], where continuous functions are used. Suppose now that Ω = C(W) as above, but assume now for simplicity that W and Δ are disjoint (with W no longer required closed). Let d be defined by df = df|_{W}, and assume that if df = 0 then f is a scalar multiple of 1. To obtain a Lipschitz seminorm on A we again just need to put a bimodule norm on Ω. The method which is closest to the usual Lipschitz norm is to specify a nowhere zero function γ on W and set

L(f) = ∥γdf∥_{∞}

(on W, allowing value +∞). In this context however, if we set ρ = γ^{−1}, it no longer makes much sense to ask that the triangle inequality hold for ρ. About the most that is reasonable is to ask that ρ, hence γ, be positive, and that γ(x, y) = γ(y, x) for (x, y) ∈ W, x ≠ y. This is a situation which has been widely studied. Entire books [Ra, RR] have been written about the problem of finding the corresponding distance between two probability measure on X, often under the heading of “the mass transportation problem”. The function ρ is then often called a “cost function”. We should clarify that when ρ is not a metric we are dealing here with mass transportation “with transshipment permitted” [RR], not the original Monge–Kantorovich [KA] mass transportation problem, which does not permit transshipment, and may well not yield a metric. When transshipment is permitted and ρ is not a metric on X, the corresponding metric on S(X) is called the Kantorovich–Rubenstein metric [KR1, KR2]. For a fascinating survey of some recent developments concerning the original Monge–Kantorovich problem see [Ev]. When X is a finite set and W is viewed as specifying edges for a graph which has X as set of vertices, the cost function ρ is naturally interpreted as assigning lengths to the edges (though we will see a quite different interpretation in Section 12). Then the metric on X coming from L_{ρ} is the usual path-length distance on the graph. There has been much study of how to compute this path-length distance efficiently for large graphs. We remark that if one prefers to have ρ defined on all of X × X one can simply set it equal to +∞ on any (x, y), x ≠ y, which is not an edge. We remark that in the context of cost functions on compact sets there may well be no non-constant functions for which the Lipschitz seminorm is finite. As one example let X be the unit interval [0, 1], and set ρ(x, y) = |x − y|^{2}. This is, in effect, because we permit transshipment — the original Monge–Kantorovich problem is quite interesting for this particular cost function, as shown in [Ev]. It is just that the minimal cost of moving one probability measure directly to another does not then give a metric on probability measures, because it may be less costly to use two or more moves. There is a variety of other bimodule norms, such as L^{p}-norms, which one can use for various differential calculi, and these give a wide variety of metrics on probability measures [Ra]. A particularly deep application of such norms, for the case of graphs, and involving explicitly Connes ideas of non-commutative metrics, appears in [Da]. (I thank Nik Weaver for bringing this paper to my attention.) Let us now discuss briefly the case in which we have A = M_{n}, a full matrix algebra. As mentioned much earlier, one natural Lip-norm on A is just L = ∥ ∥^{∼}. Now A′ can be identified by means of the normalized trace, τ, with A itself, but equipped with the trace-norm. Then A′^{0}, as in our earlier notation, consists of the matrices with trace 0. Of course, S(A) is identified with the positive matrices of normalized trace 1. With this identification, we have

ρ_{L}(μ, ν) = trace(|μ − ν|).

This is exactly one of the metrics listed (with references) in the introduction to [ZS]. Another one listed there uses the Hilbert–Schmidt norm instead of the trace norm. Listed also is a variety of other metrics on S(M_{n}) which have appeared in various applications. But I have not checked whether they come from Lip-norms. There has also been much study of the differential geometry of S(M_{n}) for a variety of Riemannian metrics, especially the “monotone metrics”, which are closely related to operator monotone functions. Two very recent articles which contain many references to previous work on this topic are [Di, S]. But the emphasis of most of this work is not on the ordinary metric which a Riemannian metric induces on S(M_{n}), but rather on the differential geometric aspects. There is also study of the volume form which is induced, and on associated probabilistic aspects. For recent related study going in the direction of non-commutative entropy see [LR].

11. Dirac operators and differential calculi

We continue our comments of the previous section, but here we focus on how Dirac operators fit into the picture. Let A be a unital ∗-algebra equipped with a C^{∗}-norm (perhaps not complete), and let π be a faithful representation of A, that is, an isometric ∗-homomorphism of A into the algebra B(H) of bounded operators on a Hilbert space H. Let D be an essentially self-adjoint, possibly unbounded, operator on H, and assume that π(a) carries the domain of D into itself for each a ∈ A, and that on this domain [D, π(a)] is a bounded operator, and so extends uniquely to a bounded operator on H. Then, following Connes, we set

L(a) = ∥[D, π(a)]∥.

As we did earlier, it is natural to require that [D, π(a)] = 0 only when a is a scalar multiple of 1. Many important examples of this situation are now known. But in general it seems difficult to ascertain whether the corresponding metric on states gives the weak-∗ topology, though this has been shown for certain examples in [Rf]. See also [W2, W3, W5], where the sets B_{t} defined at the beginning of Section 3 are shown to be totally bounded, in fact compact, for various examples. We do not deal with this question here, but rather try to relate the bimodule picture to the Dirac picture. One direction is apparent. We view B(H) as an A-A-bimodule by setting

aT b = π(a)T π(b).

Then, although D is only affiliated with B(H), conceptually we use the inner derivation which D defines, so that

da = Dπ(a) − π(a)D = [D, π(a)].

(This, of course, is the starting point for Connes’ non-commutative differential calculus [C2].) We then note that the operator norm on B(H) is an A-A-bimodule norm, and so upon setting

L(a) = ∥[D, π(a)]∥

we obtain a Lipschitz norm, which we showed to be lower semicontinuous in Proposition 3.8.

But suppose we are given instead some first order differential calculus (Ω, d) and a bimodule norm on Ω so that we obtain the corresponding Lipschitz norm L. Can we also obtain L from a Dirac operator? For this to be possible we must have L(a^{∗}) = L(a), and L must be lower semicontinuous. As mentioned earlier, L must also fit into a family of “matrix Lipschitz seminorms”. These conditions are probably not enough in general, though I have not tried to find a counterexample. But the following superficial comments help to give some perspective. (In most of the considerations which follow the algebra structure on A is only used in order to get the Leibniz inequality. Thus much of what follows actually works for order-unit spaces.) We saw in Proposition 10.3 that we can extend (Ω, d) to obtain an inner firstorder calculus. In analogy with this idea, suppose that we can realize Ω as a subspace of B(H) for some Hilbert space H, in such a way that the norm on Ω is the operator norm, and the bimodule structure is given by two ∗- representations, π_{1} and π_{2}, of A on H, so that

aωb = π_{1}(a)ωπ_{2}(b)

for a, b ∈ A and ω ∈ Ω. Suppose further that there is a possibly-unbounded essentially self-adjoint operator, D_{0}, on H, such that π_{1}(a) and π_{2}(a) carry the domain of D_{0} into itself, and such that

da = D_{0}π_{2}(a) − π_{1}(a)D_{0},

which in particular must be a bounded operator. Set L(a) = ∥da∥. This is not exactly the Dirac operator setting, but it is not difficult to convert it into that setting. To arrange matters so that we have only one representation, we let π = π_{1} ⊕ π_{2} on H ⊕ H and set

D_{1} = ( \begin{matrix}0 & D_{0} \\ 0 & 0 \end{matrix} ) .

Then we find that

L(a) = ∥[D_{1}, π(a)]∥.

But of course D_{1} is not self-adjoint. We fix this in the traditional way by again doubling the Hilbert space, with representation π ⊕ π of A, and setting

D = ( \begin{matrix}0 & D_{1}^{∗} \\ D_{1} & 0 \end{matrix} ) .

The corresponding Lipschitz norm is L(a)∨L(a^{∗}), but from the self-adjointness of D one can check that we actually get back L. Anyway, we are left with

11.1 Question. For an order-unit space A, or a ∗-algebra A with C^{∗}-norm, how does one characterize those Lip-norms on A which come from the Dirac operator construction?

Even for finite-dimensional commutative C^{∗}-algebras it is not clear to me what the answer is. As mentioned earlier, a Dirac operator also gives seminorms on all of the matrix algebras over A, so that one can speak of this family as a “matrix Lipschitz norm”, in the spirit of [Ef]. Thus a related problem is to characterize these structures. Of course a given metric on S(A) may come from several fairly different Dirac operators. For example, suppose that we have a compact space X, and a closed neighborhood W of the diagonal Δ of X × X, together with a cost function ρ on W, just as in the previous section. As discussed there, we can use ρ together with the first-order calculus determined by W to define a Lipschitz norm on C(X). (Further hypotheses are needed for it to be a Lip-norm on a dense subalgebra of C(X).) Then by the procedure discussed earlier in the present section we can pass to a Dirac operator. But that procedure enlarged the Hilbert space because a first-order differential calculus usually involves two representations rather than one. We will now show that there is an alternative method which does not enlarge the Hilbert space. This is a mild generalization of my lecture comments for metric spaces mentioned earlier, whose details are indicated on page 274 of [W2]. As earlier, let m be a measure on X of full support, and consider m × m on W ∖ Δ. Form the Hilbert space H = L^{2}(W ∖ Δ, m × m). We consider only the representation π of A = C(X) on H defined by

(π_{f}ξ)(x, y) = f(x)ξ(x, y).

(This is, of course, essentially the left action on the bimodule for W.) Define an operator, F, on H by the flip

(F ξ)(x, y) = ξ(y, x).

Because we are using a product measure, the operator F is self-adjoint and unitary. Define an (unbounded) positive operator, P, on H by

(P ξ)(x, y) = ξ(x, y)/ρ(x, y).

Because we assume that ρ(x, y) = ρ(y, x) for all (x, y) ∈ W, the operators F and P commute. We define the Dirac operator by

D = P F,

so that F is the phase of D and P = |D|. Informal calculation shows that for any f ∈ C(X) we have

([D, π_{f}]ξ)(x, y) = ((f(y) − f(x))/ρ(x, y))ξ(y, x),

so that

L(f) = ∥[D, π_{f}]∥ = sup{|f(y) − f(x)|/ρ(x, y) : (x, y) ∈ W }.

Of course, further hypotheses must be placed on ρ in order for this to give a Lip-norm. But the right-hand side of the above equality is the usual definition of a Lipschitz norm in this situation, especially in contexts such as graph theory. It will coincide with what one obtains in the corresponding bimodule approach. Notice that the resulting distance between two points x, y ∈ X can easily be strictly smaller than ρ(x, y) (if (x, y) happens to be in W). For an interesting alternative (but closely related) method of obtaining the usual distance on a graph (including infinite graphs) from a cost function, by means of Dirac operators, see theorem 7.2 of [Da]. Furthermore, in [Da] other very interesting and quite different Dirac operators associated to cost functions on graphs are discussed in some detail, and used to obtain improved estimates for heat kernels on graphs. They can be described in terms of firstorder differential calculi and Laplace operators along much the same lines as we used in Section 10. Much of this is explicit in [Da], and we will not elaborate on it here. We should mention here that very interesting examples of Dirac operators associated with non-commutative variants of sub-Riemannian manifolds appear in the second example following axiom 4′ of [C3], and in [W5].

12. Resistance distance

We conclude with an appealing class of examples which do not fit into the previous framework of differential calculi, and for which the Lip-norm does not satisfy the Leibniz identity. These examples come from graphs with “cost functions” on the edges, but now the graph is interpreted as an electrical circuit with resistances on the edges, whose values are given by the cost function. These examples have been extensively studied [DS, Kl, KlR, KZ], but I have not seen earlier mention of the corresponding metric on probability measures which we will define here. It is not clear to me whether this metric is more than a curiosity. All of the discussion here can be carried out for infinite graphs, along the lines discussed extensively in [DS], but for simplicity we only discuss finite graphs here. The examples also have a fine alternative interpretation in terms of random walks [DS]. Our term “resistance distance” is taken from the title of [KlR]. The set-up, as indicated above, is a finite graph with set X of vertices, together with strictly positive real numbers r_{xy} = r_{yx} assigned to each (undirected) edge. We interpret these numbers as resistances. We assume throughout that the graph is connected. Given x, y ∈ X, x ≠ y, we can imagine putting a voltage difference across x and y, adjusted so that one unit of current flows in at x and out at y. Then Ohm’s law says that the “effective resistance” is equal to the required voltage difference. We denote this effective resistance by ρ(x, y). It is, in fact, a metric on X. The only reference I know for this is [KlR, K, KZ], but my friends in probability theory tell me that within the context of random walks rather than resistances this is well-known, even if no reference comes to mind.

Suppose now that μ and ν are general probability measures on X. Although it does not seem so intuitively obvious, we will see shortly that we can establish voltages on the points of X such that unit total current flows into the circuit, with the amount flowing in at each point x given by μ_{x}, while unit total current flows out of the circuit, with the amount at each point given by ν (with the evident interpretation when the supports of μ and ν are not disjoint). For the analysis of this situation it is useful to define a function, c, on the edges, by c_{xy} = 1/r_{xy}. This is commonly called the “conductance”. It is convenient to extend c to all of X × X by setting c_{xy} = 0 if (x, y) is not an edge (or if y = x). Let f ∈ C(X), interpreted as voltages applied to the points of X. We let df be defined as earlier for the universal calculus (or for the calculus corresponding to the edges). We let ∇f denote the resulting flow inside the circuit. By Ohm’s law the flow (before electrons were discovered) from x to y is given by

(∇f)(x, y) = (f(x) − f(y))c_{xy} = −c(df),

where by c(df) we mean the pointwise product of functions. Note that ∇f is a function on directed edges, with

(∇f)(x, y) = −(∇f)(y, x)

(and value 0 if (x, y) is not an edge). Suppose now that ω is any function on directed edges such that ω(x, y) = −ω(y, x). We interpret ω(x, y) as giving the magnitude of a current from x to y. (To be more realistic we should require 0 circulation, but we will have no need to impose this requirement.) To sustain this current, we will in general have to insert (or extract) current at various vertices. We let div(ω)(x) denote the current which must be inserted at x. By Kirchhoff’s laws we have

div(ω)(x) = ∑_{y} ω(x, y).

Note that because ω(x, y) = −ω(y, x), we will have

∑_{x} div(ω)(x) = 0,

which accords with the fact that the total amount of current inserted must be 0. Suppose now that f ∈ C(X) and that we set ω = ∇f. We see from above that the currents which must be inserted to sustain the voltages given by f must be

div(∇f),

which we denote by Δf. To accord with our earlier notation, we let A′^{0} denote the signed measures, λ, on X for which ⟨1, λ⟩ = 0. The discussion of the previous paragraph can be interpreted as saying that Δf ∈ A′^{0}.

Suppose now that we are given λ ∈ A′^{0}. Can we find f such that Δf = λ? Note that since Δ1 = 0, we know that f will not be unique, but rather that, as usual with potential functions, we can expect f to be unique only up to a constant function. To proceed further we must more carefully analyze the operator Δ in the traditional way [DS, K]. For f ∈ C(X) we have

(Δf)(x) = ∑_{y}(∇f)(x, y)

= ∑_{y}(f(x) − f(y))c_{xy} = f(x) ∑_{y} c_{xy} − ∑_{y} f(y)c_{xy}.

Let D denote the diagonal matrix with diagonal entries

D_{xx} = ∑_{y} c_{xy}.

If we view f as a column vector, we see that

Δf = (D − C)f.

From the Peron–Frobenius theorem and the fact that our graph is connected, it follows that the kernel of Δ consists exactly of the constant functions. If we permit ourselves to confuse vector spaces a bit, we see that Δ is self-adjoint with respect to the standard inner-product on column vectors. Thus it carries the orthogonal complement, H, of the constant functions into itself, and it is invertible on H. Consequently, for every λ ∈ A′^{0} we can find a unique f ∈ H such that Δf = λ. We will write this as f = Δ^{−1}λ, where we view Δ as restricted to H so that it is invertible there. Suppose now that x and y are fixed points of X, and that λ = δ_{x} − δ_{y}, where δ_{x} denotes the δ-measure at x. Thus we are inserting one unit of current at x and extracting it at y. Let f = Δ^{−1}λ. According to our earlier comments, the effective resistance from x to y, ρ(x, y), is given by f(x) − f(y) = (Δ^{−1}λ)(x) − (Δ^{−1}λ)(y). It is now easy to see why ρ is a metric, along the lines given in [KlR]. If z is any other point of X, let

g = Δ^{−1}(δ_{x} − δ_{z}), h = Δ^{−1}(δ_{z} − δ_{y}).

Clearly f = g + h, so

ρ(x, y) = g(x) − g(y) + h(x) − h(y).

But simple considerations show that g must take its maximum and minimum values at x and z, so that

g(x) − g(y) ≤ g(x) − g(z) = ρ(x, z).

Similarly h(x) − h(y) ≤ ρ(x, z). The triangle inequality for ρ follows. But we are interested more generally in the effective resistance between μ and ν where μ and ν are arbitrary probability measures, and it is not even clear how this should be defined. (It does not seem natural just to use the Monge– Kantorovich metric from ρ.) In view of our earlier considerations we should form λ = μ − ν, and so we need an appropriate norm on A′^{0}, and this should be the dual norm of a Lip-norm, say L, on C(X), probably defined by means of a norm on Ω^{u}. The dual norm, L′, should be such that if λ = δ_{x} − δ_{y}, then L′(λ) = (Δ^{−1}λ)(x) − (Δ^{−1}λ)(y). But as remarked above, Δ^{−1}λ takes its maximum and minimum values at x and y. Thus a norm which will meet this requirement is

L′(λ) = 2∥Δ^{−1}λ∥_{∞}^{∼},

where ∥ ∥_{∞}^{∼} is as defined in Section 1. To find L on C(X) we use the selfadjointness of Δ to calculate, for g ∈ C(X) and any λ ∈ A′^{0},

⟨g, λ⟩ = ⟨g, ΔΔ^{−1}λ⟩ = ⟨Δg, Δ^{−1}λ⟩.

The supremum over λ such that 2∥Δ^{−1}λ∥_{∞}^{∼} ≤ 1 is the same as the supremum of

⟨(1/2)Δg, h⟩

over h such that ∥h̃∥_{∞}^{∼} ≤ 1. But we saw earlier that this gives just the restriction to A′^{0} of the dual norm for ∥ ∥_{∞} on C(X), which is the L^{1}-norm. Thus we see that we must set

L(g) = (1/2)∥Δg∥_{1} = (1/2) ∑_{x} |(Δg)(x)|

= (1/2) ∑_{x} |∑_{y}(g(x) − g(y))c_{xy}| = (1/2) ∑_{x} |∑_{y} dg(x, y)c_{xy}| .

This is certainly rather different from the usual Lip-norms for metrics on finite sets. The above expression suggests that we define a seminorm, N, on Ω^{u} by

N(ω) = (1/2) ∑_{x} |∑_{y} ω(x, y)c_{xy}| ,

so that we have

L(g) = N(dg).

Reversal of the earlier calculation shows that the dual norm is the L′ considered above, so that we obtain the desired ρ(μ, ν). However N will not usually be a bimodule norm, so that we are not fully in the context of the previous sections, and L need not satisfy the Leibniz inequality. I must admit that I see no particularly natural interpretation for L(g), nor for ρ(μ, ν), even if we call the latter “effective resistance”. If g were interpreted as giving voltages on X, then L(g) would be half the sum of the absolute values of the currents inserted or extracted from the circuit, and thus exactly the sum of the currents inserted into the circuit (disregarding the currents extracted). But I do not see why it is natural to give g such an interpretation as voltages. If one goes back to the effective resistance between two points, then it is easily seen that this is equal to the energy dissipated by the circuit when one unit of current is inserted. This suggests using the dissipated energy in the more general case of arbitrary probability measures μ and ν. But the energy dissipated along any edge varies as the square of the current, and one can see by examples that this causes the triangle inequality to fail. One does obtain a metric if one uses the square-root of the dissipated energy, but this does not give the correct value for the effective resistance between two points. These possibilities are not far from the Lipschitz norm used right after lemma 4.1 of [Da] to define the metric denoted there by d_{3}. This Lipschitz norm can be interpreted as the supremum over the points x of X of the square roots of the energy dissipations in all the edges beginning at x. Perhaps the discussion of Dirichlet spaces given in section 6 of [W6], or the “twisted bimodule structure” and corresponding differential discussed beginning on page 149 of [Me] in connection with Hudson’s treatment of discrete flows and stochastic differential equations, could be used to shed more light on this. Or perhaps some of the stopping rules or mixing times considered for Markov chains, as discussed in [LW], are relevant. Finally, we remark that it would be interesting to study resistance distance in the continuous case, for example for thin plates of resistance metal of various shapes.

References

[Al] Alfsen, E. M., Compact convex sets and boundary integrals, Springer-Verlag, Berlin, New York, 1971.

[Ar] Arveson, W. B., The harmonic analysis of automorphism groups, MR 84m:46085, Proc. Sympos. Pure Math., vol. 38, Amer. Math. Soc., Providence, RI, 1982, pp. 199–269.

[Ba] Baaj, S., Multiplicateurs non bornés, Thèse 3eme cycle, Université Paris VI (1980).

[BCD] Bade, W. G., Curtis, P. C., and Dales, H. G., Amenability and weak amenability for Beurling and Lipschitz algebras, MR 88f:46098, Proc. London Math. Soc. (3) 55, no. 2 (1987), 359–377.

[Bh] Bhatia, R., Matrix Analysis, MR 98i:15003, Springer-Verlag, New York, New York, 1997.

[BC] Blackadar, B. and Cuntz, J., Differential Banach algebra norms and smooth subalgebras of C^{∗}-algebras, J. Operator theory 26 (1991), 255– 282.

[C1] Connes, A., Compact metric spaces, Fredholm modules and hyperfiniteness, Ergodic Theory and Dynamical Systems 9 (1989), 207–220.

[C2] Connes, A., Noncommutative Geometry, Academic Press, San Diego, 1994.

[C3] Connes, A., Gravity coupled with matter and the foundation of non commutative geometry, hep-th/9603053., Comm. Math. Phys. 182 (1996), 155–176.

[CM] Connes, A. and Moscovici, H., Cyclic cohomology, the Novikov conjecture and hyperbolic groups, MR 92a:58137, Topology 29, no. 3 (1990), 345–388.

[Cw] Conway, J. B., A Course in Functional Analysis, 2nd edition, Springer Verlag, New York, 1990.

[Da] Davies, E. B., Analysis on graphs and noncommutative geometry, MR 93m:58110, J. Funct. Anal. 111, no. 2 (1993), 398–430.

[Di] Dittmann, J., On the curvature of monotone metrics and a conjecture concerning the Kubo-Mori metric, quant-ph/9906009.

[DS] Doyle, P. G. and Snell, J. L., Random Walks and Electric Networks, MR 89a:94023, Mathematical Association of America, Washington, DC, 1984.

[Ef] Effros, E. G., Advances in quantized functional analysis, MR 89e:46064, Proceedings of the International Congress of Mathematicians, Vol. 1, 2 (Berkeley, Calif., 1986), Amer. Math. Soc., Providence, RI, 1987, pp. 906–916.

[Ev] Evans, L. C., Partial differential equations and Monge–Kantorovich mass transfer, available at: http://berkeley.math.edu/ evans/, preprint.

[H] Hanin, L. G., Kantorovich-Rubinstein norm and its application in the theory of Lipschitz spaces, MR 92i:46026, Proc. Amer. Math. Soc. 115, no. 2 (1992), 345–352.

[J] Johnson, B. E., Amenability and weak amenability for Beurling and Lipschitz algebras, MR 88f:46098, Proc. London Math. Soc. (3) 55, no. 2 (1987), 359–377.

[K1] Kadison, R. V., A representation theory for commutative topological algebra, MR 13, 360b, Mem. Amer. Math. Soc. 1951, no. 7 (1951), 39 pp..

[K2] Kadison, R. V., Unitary invariants for representations of operator algebras, MR 19, 665e, Ann. of Math. (2) 66 (1957), 304–379.

[K3] Kadison, R. V., Transformations of states in operator theory and dynamics, MR 29 #6328, Topology 3, suppl. 2 (1965), 177–198.

[KA] Kantorovič, L. V. and Akilov, G. P., Functional Analysis (Russian), MR 86m:46001, Third edition, “Nauka”, Moscow, 1984.

[KR1] Kantorovič L. V. and Rubinšteĭn, G. Š., On a functional space and certain extremum problems, MR 20 #1219, Dokl. Akad. Nauk SSSR (N.S.) 115 (1957), 1058–1061.

[KR2] Kantorovič L. V. and Rubinšteĭn, G. Š., On a space of completely additive functions, MR 21 #808, Vestnik Leningrad. Univ. 13, no. 7 (1958), 52–59.

[Kl] Klein, D.J., Graph geometry, graph metrics, and Wiener, MR 98e:05103, Match No. 35 (1997), 7–27.

[KlR] Klein, D. J. and Randić, M., Resistance distance, MR 94d:94041, J. Math. Chem. 12, no. 1-4 (1993), 81–95.

[KZ] Klein, D. J. and Zhu, H.-Y., Distances and volumina for graphs, MR 99f:05032, J. Math. Chem. 23, no. 1-2 (1998), 179–195.

[LR] Lesniewski, A. and Ruskai, M. B., Monotone Riemannian metrics and relative entropy on non-commutative probability spaces, mathph/9808016..

[LW] Lovász L. and Winkler, P., Mixing times, Microsurveys in Discrete Probability (Princeton, NJ), Amer. Math. Soc., Providence, RI, 1997, pp. 85–133.

[Me] Meyer, P.-A., Quantum probability for probabilists, MR 94k:81152, Lecture Notes in Math., 1538, Springer-Verlag, Berlin (1993).

[MW] Moscovici, H. and Wu, F.-B., Index theory without symbols, MR 96g:58184, Contemp. Math. 167 (1994), Amer. Math. Soc., Providence, RI, 304–351.

[P] Pavlović, B., Defining metric spaces via operators from unital C^{∗}- algebras, Pacific J. Math. 186, no. 2 (1998), 285–313.

[Ra] Rachev, S. T., Probability Metrics and the Stability of Stochastic Models, John Wiley and Sons, 1991.

[RR] Rachev, S. T. and Rüschendorf, L., Mass Transportation Problems. Vol. I, Theory, Springer-Verlag, New York, New York, 1998.

[Rf] Rieffel, M. A., Metrics on states from actions of compact groups, math.OA/9807084, Doc. Math. 3 (1998), 215–229.

[Ru] Rudin, W., Functional Analysis, Second edition, McGraw-Hill, Inc., New York, New York, 1991.

[S] Slater, P. B., A priori probabilities — based on volume elements of monotone metrics — of quantum disentanglements, quant-ph/9810026, J. Phys. A32 (1999), 5261.

[W1] Weaver, N., Lipschitz Algebras, World Scientific, Singapore, 1999.

[W2] Weaver, N., Lipschitz algebras and derivations of von Neumann algebras, MR 97f:46081, J. Funct. Anal. 139, no. 2 (1996), 261–300.

[W3] Weaver, N., α-Lipschitz algebras on the noncommutative torus, J. Operator Theory 39 (1998), 123–138.

[W4] Weaver, N., Operator spaces and noncommutative metrics, preprint.

[W5] Weaver, N., Sub-Riemannian metrics for quantum Heisenberg manifolds, math.OA/9801014.

[W6] Weaver, N., Lipschitz algebras and derivations, II: Exterior differentiation, math.FA/9807096.

[Wo] Woronowicz, S. L., Unbounded elements affiliated with C^{∗}-algebras and noncompact quantum groups, MR 92b:46117, Comm. Math. Phys. 136, no. 2 (1991), 399–432.

[ZS] Zyczkowski, K. and W. Słomczyński, W., The Monge distance between quantum states, quant-ph/9711011, J. Phys. A 31, no. 45 (1998), 9095– 9104.

Marc A. Rieffel Department of Mathematics University of California Berkeley, California 94720-3840 USA rieffel@math.berkeley.edu